### v1

# Learning Object Compliance via Young’s Modulus from Single Grasps with Camera-Based Tactile Sensors

Abstract
Compliance is a useful parametrization of tactile information that humans often utilize in manipulation tasks. It can be used to inform low-level contact-rich actions or characterize objects at a high-level. In robotic manipulation, existing approaches to estimate compliance have struggled to generalize across object shape and material. Using camera-based tactile sensors, we present a novel approach to parametrize compliance through Young’s modulus $E$. We evaluate our method over a novel dataset of 285 common objects, including a wide array of shapes and materials with Young’s moduli ranging from 5.0 kPa to 250 GPa. Data is collected over automated parallel grasps of each object. Combining analytical and data-driven approaches, we develop a hybrid system using a multi-tower neural network to analyze a sequence of tactile images from grasping. This system is shown to estimate the Young’s modulus of unseen objects within an order of magnitude at 74.2% accuracy across our dataset. This is a drastic improvement over a purely analytical baseline, which exhibits only 28.9% accuracy. Importantly, this estimation system performs irrespective of object geometry and demonstrates robustness across object materials. Thus, it could be applied in a general robotic manipulation setting to characterize unknown objects and inform decision-making, for instance to sort produce by ripeness.
Keywords: Robotic Manipulation, Tactile Sensing, Compliance
1 Introduction
Humans use tactile information to characterize and interact with objects [1]. Through touch, we sense dynamic properties such as friction and compliance [2]. In the case of compliance, we detect discrepancies between soft and hard objects through fingertip pressure and deformation [3].
Compliance has been established as a useful parameter in robotics to classify objects and inform safe grasping policies [4]. It can be directly utilized for a wide variety of biomedical applications [5, 6, 7] or sorting produce by ripeness [8, 9, 10]. Compliance has also been shown to be useful for manipulation through haptic teleoperation systems [11]. Moreover, a reliable measure of compliance could greatly benefit an intelligent, LLM-backed planning architecture [12, 13].
Tactile sensors deployed on robotic end-effectors offer a similar capability to detect dynamic properties as human fingers [14]. High-resolution, camera-based tactile sensors like GelSight have demonstrated strong performance in perceiving properties such as friction and compliance [15, 16]. Incorporating these sensors into robotic systems has proven to be beneficial for executing dynamic tasks [17, 18], including manipulating deformable objects such as cloth or cables [19, 20].
Even with the help of high-resolution tactile sensing, compliance is particularly difficult to discern in unstructured environments. An object’s deformation in response to loading depends both on the object’s shape and its compliance. Without known geometry and pose tracking, analytical methods cannot be utilized to rigorously determine compliance. Instead, learning-based implementations have been developed to estimate compliance independent of shape [16, 6, 12, 13]. These methods demonstrate relative success in estimating compliance across geometries. However, there remains a challenge of universally describing compliance for all material types.
To better generalize across both shapes and materials, we present a novel compliance estimation system which incorporates analytical modeling alongside learning. Specifically, we use a neural network to combine CNN encoded tactile images with estimates produced by analytical contact models. This hybrid approach allows us to better generalize across materials given an analytical backbone while using learning techniques to account for complex geometries. We validate our approach with a novel dataset of 285 common objects which includes a wide variety of shapes and materials. Our estimation system is shown to perform well across this large, diverse dataset.
Young’s modulus $E$ is a property that defines a material’s resistance to deformation in the elastic regime [21]. We choose to parametrize compliance using Young’s modulus because it is universally defined for all materials and is commonly utilized in well-established mechanical models. For these reasons, it is a natural choice for quantifying compliance in robotic manipulation.
Given an increased ability to generalize across object shape and material type, our compliance estimation architecture could be applied in general robotic manipulation scenarios to characterize unknown objects. More acutely, it could be applied to tasks such as produce sorting to identify ripeness of fruits and vegetables [13, 8, 10].
Figure 1: Young’s modulus estimation system. Tactile images from parallel grasps of various objects are fed into learned models alongside analytical estimates to generate a hybrid estimate of Young’s modulus. Tactile images are size 35mm by 25mm.
2 Related Work
2.1 GelSight Sensors & Contact Modeling
Camera-based tactile sensors, such as GelSight sensors, are designed to track contact geometry in high-resolution using a camera placed behind an elastomer gel pad [15, 22, 23, 24, 25, 26]. They are highly effective in texture recognition [27, 28, 29], shape and pose estimation [30, 31], and measuring contact force [15] and friction [32]. However, reliably modeling contact with these soft sensors remains challenging. Simulators have been developed to generate tactile images and deformed geometry of sensors using pre-computed finite element method (FEM) solutions [33]. Finite element analysis has proven a viable way to model soft contact on hardware and used to estimate forces from tactile data [34, 35]. Though modeling with FEM is accurate, it requires a precise map of object geometry and solutions must be pre-computed offline.
Alternative approaches have been developed using Hertzian mechanics to model soft contact in robotic grasping [36, 37]. This closed-form method offers improved computational efficiency over FEM and simulated methods. Still, it requires assumptions on contact geometry, namely that geometry should be axisymmetric. Hertzian contact models have been utilized for force estimation [38, 6] and surface curvature detection from tactile data [39]. Since they are closed-form, these constitutive models can more easily be inverted to determine object compliance.
2.2 Compliance Estimation
Estimating the compliance of a contacted object remains a complex problem in robotic manipulation, particularly without assumptions on object shape or material type. Analytical approaches to compliance estimation often lack robustness to variations in geometry. A traditional approach to model compliance is to apply Hooke’s Law using measurements of grasping force and positions [4]. Similar models have been foundational in developing low-level grasping control policies [40, 41]. This method can provide an effective stiffness of grasping. However, it requires assumptions on object shapes and may struggle when handling complex geometries without a known map of each grasped object. By incorporating tactile sensing, it is possible to create higher fidelity compliance estimation models [42, 43, 44]. These methods have a stronger ability to sense and generalize across contact geometry but still struggle to account for complex contact geometries.
Recent learning-based approaches to compliance estimation have demonstrated the ability to generalize across geometry but struggle across a wide-range of material types. Work has been done to create a tactile-informed model to estimate the Shore 00 hardness of an object in contact with a GelSight sensor. The architecture predicts hardness directly from tactile images without any mechanical modeling [45, 16]. Shore 00 hardness is a measure designed to quantify the compliance of soft rubbers [46]. Thus, this method is trained using only rubber objects. It is fundamentally restricted to predicting on this scale and cannot express differences in compliance between more rigid objects. LLM-backed methods have been developed to classify tactile images with binary semantic descriptions of ‘soft’ or ‘hard’ [12, 13]. These large-data methods are promising but still have not been used to assess compliance on a continuous scale across material types. Other data-driven implementations have been developed using tactile sensing to directly measure the ripeness of produce [8, 9, 10]. As before, these methods are restricted by their dataset and labeling method. In all cases, labeling choices prevent models from incorporating analytical modeling alongside learning.
3 Methods
3.1 Overview
We develop a hybrid system fusing analytical models alongside learning to estimate the Young’s modulus of general unknown objects from a single parallel grasp. Through analytical modeling, we can create a well-founded, preliminary estimate for Young’s modulus. By incorporating learning, we can compensate for assumptions taken by analytical models to better generalize across contact geometry. A large, diverse dataset of 285 common objects is collected to train and evaluate our method. Our approach of combining analytical models with data-driven methods is partially inspired by residual physics architectures [47] and physics-informed tactile models [48].
Our hybrid estimation architecture is displayed in Fig. 2. Tactile images and measurements of force $F$ and gripper width $w$ are fed into Hertzian and elastic analytical models to fit respective estimates of Young’s modulus $\hat{E}_{\text{elastic}},\hat{E}_{\text{hertz}}$. In parallel, a selected number $N_{\text{frames}}$ of tactile images are fed into a data-driven model. Features extracted from images are concatenated with grasp measurements into a large fully-connected decoder. This multi-tower architecture is similar to other grasp-based learning architectures [49]. Output features from our learned model are sent to a smaller decoder alongside analytical estimates to produce a final hybrid estimate of Young’s modulus $\hat{E}$.
Figure 2: Young’s modulus estimation architecture. Tactile images, depth, and measured grasping force $F$ and width $w$ are sent into analytical and learned models. Estimates from analytical models are fed into a fully-connected decoder alongside learned grasp features.
Code for the project is publicly available on GitHub. Collected data for multiple grasps of all objects with labeled Young’s moduli and Shore hardness is publicly available on HuggingFace.
3.2 Analytical Contact Models
Simple Elasticity Parallel grasping can be modeled as uniaxial loading using an adaptation of Hooke’s Law. A diagram of parallel grasping is provided in Fig. 3. At every instance $t$, force $F$ is applied to the grasped object, with gripper at width $w$. Traditional methods have used these measurements alone to estimate the compliance or stiffness of grasped objects [4].
With camera-based tactile sensing, we can increase the fidelity of the model by measuring contact area $A$ and depth $d$ [15]. Contact area $A$ is determined by masking depth images with constant threshold of 0.1mm. This threshold was chosen empirically to cancel out potential noise. Without tactile sensing, we could not measure depth ($d(t)=0$) and would need to assume constant contact area ($A(t)=A_{\text{sensor}}$). We use measurements to calculate contact stress $\sigma$ and object strain $\epsilon$, following Eq. 1. We define $t=0$ to be the moment of first contact, where $F(t=0)$ must be greater than an empirically-defined threshold.
$\sigma(t)=\frac{F(t)}{A(t)}\;,\quad\quad\epsilon(t)=\frac{\Delta w(t)+2\Delta d(t)}{w(0)+2d(0)}=\frac{(w(t)+2d(t))-(w(0)+2d(0))}{w(0)+2d(0)}$ (1)
In the elastic regime, stress $\sigma$ and $\epsilon$ are related by Young’s modulus [21]. Using measurements across grasping time $t$, we fit an estimate for Young’s modulus $\hat{E}_{\text{elastic}}$ following Eq. 2.
$\boldsymbol{\sigma}=\hat{E}_{\text{elastic}}\boldsymbol{\epsilon}$ (2)
Despite its simplicity, this method does not properly consider the geometry of unknown objects. It implicitly assumes objects have a consistent profile along the grasping axis. Furthermore, it does not take advantage of precise information about the deformation of contact geometry that is apparent through high-resolution tactile data.
Figure 3: Simple mechanics of parallel grasping. Spherical object of radius $R$ is grasped with soft GelSight sensors. Normal force $F$ is applied at width $w$. Penetration depth $d$ and contact radius $a$ can be measured with tactile sensors. Timestep $t=0$ is the moment of first contact over the grasp.
Hertzian Contact Hertzian contact theory examines the deformation of elastic bodies through contact [50]. In a simple case, contact is assumed to be without friction or adhesion [51, 52]. Unlike the previously-described method, Hertzian mechanics more acutely analyzes geometry through contact. It has been used to model soft contact in robotics [36, 39], including GelSight-like tactile sensors [23, 6]. Precise depth sensing allows us to track surface deformation and apply this model.
GelSight sensors are made with highly elastic silicone rubber [22]. Unless the grasped object is fully rigid, we expect both bodies to deform through contact. This implies the maximum measured penetration depth $d$ will be less than the relative displacement between the finger and object $d<\frac{\Delta w}{2}$. This phenomenon is depicted in Fig. 3. An aggregate modulus $E^{*}$ describes the bodies’ combined resistance to deformation in Eq. 3, which depends on the Poisson’s ratio $\nu$ of each body.
$\frac{1}{E^{*}}=\frac{1-\nu_{\text{sensor}}^{2}}{E_{\text{sensor}}}+\frac{1-\nu_{\text{obj}}^{2}}{E_{\text{obj}}}$ (3)
Hertzian solutions have been derived for a wide range of contact geometries [53]. Since we are analyzing contact with an unknown object whose shape is not precisely known, we model all contact geometry as spherical. This is a reasonable simplification for convex objects, but may produce bounded errors in estimates over our geometrically diverse dataset. Our neural network will learn to compensate for this intrinsic and repeatable geometric modeling error.
Since we assume contact profiles are spherical, we must calculate contact radius $a$. This is the radius needed for a circle’s contact area to match the observed contact area $A$. Additionally, we can approximate the apparent radius of the unknown object $R$ through observed contact area and gripper displacement [53]. Using these parameters and the method of dimensionality reduction (MDR), we can calculate the sensor’s surface deformation in response to elastic loading [53, 54]. This relates force $F$ and depth $d$ by $E^{*}$. This relation shows that a harder sphere will plunge deeper into the contact plane, assuming constant force. More detailed derivation for this expression and discussion about the MDR is provided in Appendix A.
$a(t)=\sqrt{\frac{A(t)}{\pi}},\quad R(t)=\frac{a^{2}(t)}{\Delta w(t)},\quad d(t)=\frac{1-\nu_{\text{sensor}}^{2}}{E_{\text{sensor}}}{\left(\frac{3{E^{*}}^{2}F(t)}{32R^{2}(t)}\right)}^{\frac{1}{3}}a(t)$ (4)
With measured data over the duration of a grasp, we apply least-squares to retrieve a best fit value $\hat{E}^{*}$. Then, we use $\hat{E}^{*}$ to compute an estimated modulus for our unknown object $\hat{E}_{\text{hertz}}$, given known properties of our sensor $E_{\text{sensor}}=0.275$ MPa, $\nu_{\text{sensor}}=0.48$. We assume constant $\nu_{\text{obj}}=0.4$ based on the Poisson’s ratio of common materials [55]. After analytical estimates are computed, linear scaling is added to account for potential constant errors across the dataset.
$\hat{E}_{\text{hertz}}=\left(1-\nu_{\text{obj}}^{2}\right){\left(\frac{1}{\hat{E}^{*}}-\frac{1-\nu_{\text{sensor}}^{2}}{E_{\text{sensor}}}\right)}^{-1}$ (5)
4 Physical Dataset
A novel physical dataset of 285 objects was gathered to train and validate our method. For comparison, our dataset is nearly 4 times larger than the YCB dataset, which includes only 77 objects [56]. Chosen objects were intended to be among things a robotic manipulator may commonly interact with in general operation. Additionally, each object was required to have approximately uniform material composition to ensure that an accurate Young’s modulus could be defined for the object and used to label data for supervised training. A subset of dataset objects is displayed in Fig. 4.
Figure 4: Physical dataset. (a) Selected objects from our physical dataset are displayed. These are common objects with a wide variety of shapes and materials. (b) Some custom objects are molded out of silicone to replicate the shape of non-rubber objects in the dataset.
To evaluate how well estimation algorithms could generalize, diversity of shape and material was prioritized in acquiring objects. Thus, we mold 20 custom objects out of silicone to replicate the exact shape of other objects in the dataset with vastly different material composition. Molded objects are displayed in Fig. 4b. Training on data collected across object replicas of different materials will help prevent overfitting to shapes and textures. Moreover, there is incentive to bias towards rubber objects because it is clear from Eq. 5 that it will be easiest to discriminate between Young’s moduli nearest to our sensor’s. The dataset’s composition is further detailed in Appendix D.
As objects were gathered, they were labeled with their Young’s moduli. For harder objects, like those made of metals and plastics, this was done by referencing an engineering database of material properties [57], as all objects are made of known materials. For softer objects, like those made of rubber, Shore hardness was measured and converted to Young’s modulus. Shore hardness is converted to Young’s modulus using Gent’s model of hardness and other established methods [46, 58]. These models provide a reliable measure of object compliance on a logarithmic scale. The dataset includes materials ranging from foam to steel, with Young’s moduli from 5.0 kPa to 250 GPa. For reference, our sensor’s Young’s modulus is 0.275 MPa.
5 Experiments
5.1 Hardware Setup
Figure 5: Hardware setup for data collection. GelSight Wedge sensors are attached to Franka Panda gripper with embedded force gauge.
GelSight Wedge sensors [22] were rigidly attached to a standard parallel robotic gripper with 3D-printed plastic fingers, as depicted in Fig. 5. One sensor is inscribed with markers for tracking surface displacement. Markers help provide information about shear forces and surface displacement, as further explained in Appendix B. By collecting data with different sensors simultaneously, we have the ability to assess the relative necessity of markers in estimating Young’s modulus. Normal force is measured through a gauge embedded in the left finger. Gripper position is directly measured from control software.
5.2 Data Collection & Training
Figure 6: Raw force and depth data. Data is collected across grasps for spheres of each material type. Data is fit to functions of the form $d\sim F^{\frac{2}{3}}$, per our Hertzian model.
Data Collection: Automated parallel grasps are executed on objects in fixed position. For each object in our dataset, a set number of grasps are recorded. Objects are re-oriented between each grasp to diversify potential antipodal grasp locations. During grasping, tactile images from GelSight sensors, forces, and gripper widths, are recorded at 30 Hz. Measurements are synchronized by shifting for latency between cameras. Grasps are executed at constant velocity up to force of 30 N. Raw data collected from grasps of spherical objects is plotted in Fig. 6. More details about processing data collected from grasps is provided in Appendix C.
Tactile and depth images at grasping force of 30 N for a set of example objects are depicted in Fig. 7. We observe that textures are more finely captured for rigid object’s with higher modulus. As expected by Eq. 4, peak depth is higher for objects of the same shape with higher modulus.
Data Preprocessing: Data recorded from grasping objects is clipped to include only the loading sequence from first contact to peak force. Misaligned grasps where loading is non-monotonic are disregarded. From here, a set number of frames $N_{\text{frames}}=3$ are sampled equidistantly in time. Multiple augmentations of samples are taken from each grasp. As depicted in Fig. 2, tactile images are sent as inputs to our multi-tower architecture alongside analytical estimates and force and width measurements. Images are randomly transformed with horizontal and vertical flipping.
Figure 7: Example tactile data. Tactile images and depth reconstruction from a GelSight Wedge sensor without markers over a set of example objects. Shown frames are all from peak pressing force of 30 N. Tactile images are size 35mm by 25mm.
Training: We train our model for 80 iterations over 4,000 grasp inputs. In total, these grasps include more than 12,000 tactile images. Output labels are normalized on a log10 scale given the range of Young’s moduli in our dataset. Final sigmoid activation function is used with an MSE loss function.
6 Experimental Results
Figure 8: Young’s modulus predictions for rubber objects only. Results for model trained on only rubber objects over 80 randomly sampled grasps of unseen objects. Predictions in the gray region are considered sufficiently accurate.
We train our neural network with preprocessed data as described in Section 5.2. Initially, we train a model over only rubber objects. This preliminary exercise is performed to evaluate the model’s ability to learn across a constrained dataset. Rubber objects are closest to the Young’s modulus of our sensor. From Eq. 5, we expect it will be easier to discern finer differences in compliance between them. Results over unseen objects for this model are plotted in Fig. 8. The model demonstrates 100% prediction accuracy within an order of magnitude of the ground truth Young’s modulus. The prediction range of this experiment is analogous to previously developed Shore hardness tactile models [16], which are designed to estimate the compliance of soft rubber objects. Shore 00 hardness values evaluated in [16] range from approximately 6 kPa to 2 MPa. Here, we include harder rubber objects as well.
We train our neural network over the entire collected dataset. This learned model is compared to analytical baselines. Predictions across the dataset for each method are plotted in Fig. 9. We observe that as the traditional Hooke’s Law method without tactile sensing fails to distinguish between all material types. Simple elastic and Hertzian analytical methods demonstrate significant improvements, but are still less than 30% accurate. Our learned model showcases dramatic improvements over all other methods.
All estimation methods had more difficulty expressing precise differences in Young’s modulus between harder objects that strayed further from the modulus of the sensor (i.e. plastic, wood, and metal objects). Additionally, we observe decreased overall accuracy from previous evaluation over the rubber subset of our dataset. These results are expected given that harder objects are approximately rigid relative to the soft sensor’s Young’s modulus. The resolution of the tactile sensor is likely not high enough to detect small differences in deformation between this range of compliances.
In Table 1, performance is compared between different estimation methods and inputs. Performance metrics are computed for grasps of both seen and unseen objects. We evaluate models on a logarithmic scale given the range of Young’s moduli in our dataset. Predictions are considered sufficiently accurate within the same order of magnitude of ground truth labels. We compute error as the average difference between prediction and ground truth on a log10 scale.
Figure 9: Young’s modulus predictions by material. Results for learned and analytical estimation methods are plotted, using predictions from 80 randomly sampled grasps per material category. Predictions in the gray region are considered sufficiently accurate.
From the ablation study, we observe that our learned model vastly reduces error and improves accuracy over analytical methods. Furthermore, we see that RGB tactile images with markers are favorable for this estimation task. However, we do not observe significant difference in results between models which consider force $F$ and width $w$ measurements. This may be due to the fact that all objects were grasped with nearly identical peak force. Moreover, force is implicitly detectable through tactile images [16]. Importantly, considering analytical estimates $\hat{E}_{\text{elastic}},\hat{E}_{\text{hertz}}$ is shown to improve learned prediction results. This demonstrates that incorporating analytical modeling improves our ability to discern compliance over a purely learning-based method. Accuracy and error metrics are computed across shape and material categories for each method in Appendix D. We observe that our model performs independent of shape and can generalize well across materials.
Inputs Seen Objects Unseen Objects
Method Depth / RGB Markers $F$ $w$ $\hat{E}_{\text{elastic}},\hat{E}_{\text{hertz}}$ Log10 Error Log10 Acc. Log10 Error Log10 Acc.
No Tactile 2.13 18.2%
Simple Elasticity Depth 1.90 27.3%
Hertzian (MDR) Depth 1.68 28.9%
RGB 0.73 76.9% 0.76 73.5%
RGB $\times$ $\times$ 0.73 78.5% 0.76 74.2%
RGB $\times$ 0.85 70.4% 0.88 68.6%
Top-10 NN RGB $\times$ $\times$ $\times$ 0.85 69.6% 0.90 67.9%
(Our Method) RGB $\times$ 0.63 80.4% 0.86 64.6%
RGB $\times$ $\times$ $\times$ 0.65 78.5% 0.85 65.0%
Depth 1.05 61.7% 0.98 63.6%
Depth $\times$ $\times$ 1.08 60.1% 1.04 59.4%
Table 1: Estimation Architecture Ablation Study
7 Conclusion
We have developed a novel system for estimating the Young’s modulus of unknown contacted objects using analytical and data-driven methods. This system can create estimates online from only a single parallel grasp. It has been shown to perform independent of object shape and demonstrates robustness across materials. It can expressively distinguish precise differences in compliance between soft objects. The system is less capable of measuring precise differences between harder, such as plastic and metal, objects which are effectively rigid relative to the soft GelSight sensor.
The estimation system introduced in this paper could be applied for use with other camera-based tactile sensors capable of tracking depth. With more diverse tactile data, the accuracy of the system could be further improved. Given an increased ability to generalize, this compliance estimation system could be deployed alongside an intelligent high-level controller to inform decision-making in robotic manipulation scenarios by characterizing the compliance of unknown objects. As one example, it could be utilized to identify the ripeness of fruits or vegetables [13, 8, 10].
Acknowledgments
The authors would like to thank Sandra Q. Liu, Jialiang (Alan) Zhao, Megha Tippur, Yuxiang Ma, and Laurence Willemet for their guidance and insight provided throughout the project’s duration.
Appendix A Method of Dimensionality Reduction (MDR)
As discussed in Section 3.2, we model contact between an unknown object and our camera-based tactile sensor as a collision between an elastic sphere and an elastic half-space using Hertzian contact theory. Now, we will walk through the derivation for a constitutive relation using the method of dimensionality reduction (MDR) [59]. A mechanical diagram of contact at time $t$ with labeled variables is provided in Fig. 10.
Figure 10: Method of dimensionality reduction (MDR) diagram. An elastic sphere is colliding with an elasic half-space at time $t$. Variables of contact are labeled. The two-dimensional radial stress field $p$ is transformed into a one-dimensional force field $q_{\text{1D}}$. This transformation can be observed through comparison of top and side views.
First, we compute the maximum pressure of contact $p_{0}$ between both bodies in Eq. 6. This is a simple expression of measured force $F$ and contact radius $a$. We can substitute $a$ to create an expression dependent on the aggregate elastic compliance of the bodies $E^{*}$ [52].
$p_{0}(t)=\frac{3F(t)}{2\pi a^{2}(t)}=\frac{1}{\pi}{\left(\frac{F(t){E^{*}}^{2}}{R^{2}}\right)}^{\frac{1}{3}}$ (6)
Using this maximum pressure $p_{0}$ at timestep $t$, we create the radial pressure field of contact $p(r,t)$. This expression comes directly from Hertzian contact theory [52].
$p(r,t)=p_{0}(t)\sqrt{1-\frac{r^{2}}{a^{2}(t)}}$ (7)
Now, we will apply the method of dimensionality reduction (MDR). This technique is used to solve the stress field for surface deformation based on the assumption that contact geometries are axisymmetric. By utilizing this method, we are able to reconstruct the deformation of the sensor surface. We begin by transforming coordinates of the pressure field into one-dimensional force density $q_{\text{1D}}$ through Eq. 8.
[…]
From here, we apply a constitutive relationship on the sensor gel pad to get a one-dimensional displacement of the gel surface $u_{\text{1D}}$. This displacement is defined in the normal direction $\hat{z}$. By transforming to a single dimension, we restrict the displacement geometry to be axisymmetric about $\hat{z}$ from the center $x=0$.
$u_{\text{1D}}(x,t)=\left(\frac{1-\nu_{\text{sensor}}^{2}}{E_{\text{sensor}}}\right)q_{\text{1D}}(x,t)$ (9)
The maximum displacement will occur at $x=0$. We define this to be equivalent to our maximum measured depth $d$.
$u_{\text{1D}}(0,t)=d(t)=\left(\frac{1-\nu_{\text{sensor}}^{2}}{E_{\text{sensor}}}\right)q_{\text{1D}}(0,t)$ (10)
Then, by plugging in $x=0$, the expression for $q_{\text{1D}}$ simplifies considerably.
[…]
We solve this integral to get Eq. 13 which expresses maximum observed surface displacement $d$ in terms of normal force $F$.
[…]
$u_{\text{1D}}(0,t)=d(t)=\frac{1-\nu_{\text{sensor}}^{2}}{E_{\text{sensor}}}{\left(\frac{3{E^{*}}^{2}F(t)}{32R^{2}(t)}\right)}^{\frac{1}{3}}a(t)$ (13)
We apply this relation for every timestep $t$ over the duration of a grasp. Using all measurements of $d$, $a$, and $F$, we can apply a linear least-squares algorithm to retrieve a best fit value $\hat{E}^{*}$. Then, this estimate is used with known mechanical properties of our sensor’s silicone gel pads to compute the Young’s modulus of the unknown grasped object $\hat{E}_{\text{hertz}}$, as shown in Eq. 5.
Without the MDR, we could not have solved directly for the deformation of the gel surface, meaning we could not create an expression relating $F$ directly to $d$ in terms of Young’s modulus. Instead, we could only derive a relation for the apparent deformation of both bodies with respect to each other. This apparent deformation is equal to the displacement between both bodies, or change in gripper width $\Delta w$ in our case. This is a classic tabulated solution for Hertzian contact between a sphere and a plane [52], written as Eq. 14.
$F(t)=2E^{*}a(t)\Delta w(t)$ (14)
Notably, this model would not make use of tactile depth $d$ and relies solely on gripper width $w$ measurements. These gripper measurements exhibit a larger observed error than tactile depth measurements, which are higher-resolution and demonstrate sub-millimeter accuracy. For these reasons, we opt to use the MDR-derived solution.
Appendix B Markers
Figure 11: Markers. Tactile images from GelSight Wedge’s with and without markers. Tactile images are size 35mm by 25mm.
Markers are small black dots painted in an array on the surface of the GelSight gel pad. As the surface of the gel deforms and stretches, these markers move. Thus, we can track their relative position to map the displacement of the gel pad surface in both shear and normal directions. Markers have been shown to be effective for measuring shear and normal forces from tactile images [15, 22]. From a broad standpoint, this means we can use markers to increase the amount of dynamic information encoded in a tactile image. Without them, we could not estimate shear force.
In our experiment, we have created one GelSight Wedge [22] with markers painted on the surface. This was done to investigate how beneficial markers may be for estimating a dynamic property of contact like Young’s modulus. The difference between tactile images with or without markers for two objects is tabulated in Fig. 11. These markers help track surface displacement, but they may occlude fine features of texture, like screw threads.
Appendix C Processing Grasp Data
We collect data over a set number of automated parallel grasps for every object in our dataset. Normal force $F$, width $w$, and maximum depth $d$ are plotted for an example grasp of a rigid sphere in Fig. 12a. Grasps are executed at nearly constant velocity to a peak force of approximately 30 N. These measurements are clipped to the loading sequence based on force thresholding. Once force passes an empirically-defined threshold of 0.75 N, we consider the object to be in-contact with the GelSight sensor and define this as timestep $t=0$. We clip data from this point until the peak recorded force is reached.
As expected in an elastic regime, we observe that force and depth measurements are high correlated from both GelSight sensors. Raw collected data normalized per grasp is plotted in Fig. 12b. Measurements are normalized according to their peak value over each respective grasp. These measurements are randomly sampled from grasps across the entire dataset. All measurements outside of 25% correlation bounds are discarded. Over 90% of measured points lie within these bounds for both sensors. We observe that there is no distinguishable difference between measurements collected from one sensor to the other.
Figure 12: Grasping data. (a) Grasp data is plotted for an example grasp of a rigid sphere. The chosen loading sequence from initial contact to peak force is highlighted in gray. (b) Normalized force and depth measurements from all collected grasps are plotted. The correlation coefficient between these measurements is computed as 0.84 and bounds are placed at 25% error mark.
Appendix D Dataset Diversity & Performance
The diversity of the collected dataset is characterized through categories of material and shape in Tables 3 and 3. As intended, the most common material of objects in the dataset is rubber, which includes 28.6% of the objects in the dataset. This is done to bias the dataset toward the Young’s modulus of our sensor’s gel pads. There are significant percentages of metal and plastic objects, making up 25.1% and 24.0% of the dataset respectively. Over 50% of objects in the dataset have irregular shapes. This is done to diversify training data in an attempt and learn Young’s modulus independent of object shape or contact profile. Still, there are substantial percentages of simple shapes for performance evaluation.
Table 2: Dataset Material Distributions
Material Percent (%)
Rubber 28.6%
Metal 25.1%
Plastic 24.0%
Wood 8.4%
Foam 6.3%
Other 7.6%
Table 3: Dataset Shape Distributions
Shape Percent (%)
Cylinder 20.2%
Sphere 11.5%
Rectangular 15.3%
Hexagonal 2.1%
Irregular 50.9%
Results of investigated methods on unseen objects across shape and material categories are compared in Fig. 13. We observe that neural network methods give the lowest log10 error and accuracy in nearly every category. Moreover, the accuracy increases and error decreases with the complexity of analytical models. Our Hertzian method performs best over spheres, which is expected given that this model assumes spherical contact geometry. Most critically, it is shown that our trained model can learn Young’s modulus independent of shape, given that prediction accuracy and error remain consistent across different object shapes. The model performs robustly across different categories of materials as well.
Figure 13: Estimation methods performance. Log10 error and accuracy are plotted for each estimation method across dataset categories of shape and material. Neural network results are from top-10 trained models evaluated on unseen objects.

### v2

@@
-Learning Object Compliance via Young’s Modulus from Single Grasps with Camera-Based Tactile Sensors
+Learning Object Compliance via Young’s Modulus from Single Grasps using Camera-Based Tactile Sensors
 Abstract
@@
 In robotic manipulation, existing approaches to estimate compliance have struggled to generalize across object shape and material.
-Using camera-based tactile sensors, we present a novel approach to parametrize compliance through Young’s modulus $E$.
-We evaluate our method over a novel dataset of 285 common objects, including a wide array of shapes and materials with Young’s moduli ranging from 5.0 kPa to 250 GPa.
+Using camera-based tactile sensors, proprioception, and force measurements, we present a novel approach to estimate object compliance as Young’s modulus $E$ from parallel grasps.
+We evaluate our method over a novel dataset of 285 common objects, which includes a wide array of shapes and materials with Young’s moduli ranging from 5.0 kPa to 250 GPa.
 Data is collected over automated parallel grasps of each object.
@@
 This system is shown to estimate the Young’s modulus of unseen objects within an order of magnitude at 74.2% accuracy across our dataset.
-This is a drastic improvement over a purely analytical baseline, which exhibits only 28.9% accuracy.
+This is a drastic improvement over purely analytical and data-driven baselines which exhibit 28.9% and 65.0% accuracy respectively.
 Importantly, this estimation system performs irrespective of object geometry and demonstrates robustness across object materials.
@@
 1 Introduction
-Humans use tactile information to characterize and interact with objects [1].
-Through touch, we sense dynamic properties such as friction and compliance [2].
-In the case of compliance, we detect discrepancies between soft and hard objects through fingertip pressure and deformation [3].
-Compliance has been established as a useful parameter in robotics to classify objects and inform safe grasping policies [4].
-It can be directly utilized for a wide variety of biomedical applications [5, 6, 7] or sorting produce by ripeness [8, 9, 10].
-Compliance has also been shown to be useful for manipulation through haptic teleoperation systems [11].
-Moreover, a reliable measure of compliance could greatly benefit an intelligent, LLM-backed planning architecture [12, 13].
-Tactile sensors deployed on robotic end-effectors offer a similar capability to detect dynamic properties as human fingers [14].
-High-resolution, camera-based tactile sensors like GelSight have demonstrated strong performance in perceiving properties such as friction and compliance [15, 16].
-Incorporating these sensors into robotic systems has proven to be beneficial for executing dynamic tasks [17, 18], including manipulating deformable objects such as cloth or cables [19, 20].
-Even with the help of high-resolution tactile sensing, compliance is particularly difficult to discern in unstructured environments.
-An object’s deformation in response to loading depends both on the object’s shape and its compliance.
-Without known geometry and pose tracking, analytical methods cannot be utilized to rigorously determine compliance.
-Instead, learning-based implementations have been developed to estimate compliance independent of shape [16, 6, 12, 13].
-These methods demonstrate relative success in estimating compliance across geometries.
-However, there remains a challenge of universally describing compliance for all material types.
-To better generalize across both shapes and materials, we present a novel compliance estimation system which incorporates analytical modeling alongside learning.
-Specifically, we use a neural network to combine CNN encoded tactile images with estimates produced by analytical contact models.
-This hybrid approach allows us to better generalize across materials given an analytical backbone while using learning techniques to account for complex geometries.
-We validate our approach with a novel dataset of 285 common objects which includes a wide variety of shapes and materials.
-Our estimation system is shown to perform well across this large, diverse dataset.
-Young’s modulus $E$ is a property that defines a material’s resistance to deformation in the elastic regime [21].
-We choose to parametrize compliance using Young’s modulus because it is universally defined for all materials and is commonly utilized in well-established mechanical models.
-For these reasons, it is a natural choice for quantifying compliance in robotic manipulation.
-Given an increased ability to generalize across object shape and material type, our compliance estimation architecture could be applied in general robotic manipulation scenarios to characterize unknown objects.
-More acutely, it could be applied to tasks such as produce sorting to identify ripeness of fruits and vegetables [13, 8, 10].
+General and robust object compliance estimation could significantly enhance robotic perception systems.
+For instance, a ripe avocado appears visually identical to an unripe avocado, yet their compliance differs greatly.
+Without the ability to accurately estimate compliance, a robot cannot effectively distinguish between such visually similar objects.
+Compliance measurement has proven valuable in multiple applications, including identifying the ripeness of fruits and vegetables [1, 2, 3, 4] and in biomedical fields [5, 6, 7].
+In the future, a universal compliance estimation architecture could be applied in assembly or recycling processes to inform decisions based on object material.
+In robotics, we seek to replicate the human ability to perceive properties like compliance through touch [8].
+Humans naturally use tactile information to characterize objects and inform interactions [9].
+In the case of compliance, we sense discrepancies between soft and hard objects through fingertip pressure and deformation [10].
+High-resolution, camera-based tactile sensors like GelSight are specifically designed to track the deformation of a robotic fingertip during contact [11, 12].
+However, even with the help of high-resolution tactile sensing, achieving human-like compliance perception remains challenging, especially in unstructured environments.
+An object’s deformation in response to loading depends on both its shape and material composition.
+Thus, without known geometry and pose tracking, analytical methods cannot be utilized to rigorously determine compliance [13].
+Alternatively, learning-based implementations have been developed to estimate compliance from tactile images alone without the use of mechanical modeling [14, 5, 15, 1].
+While these methods have demonstrated relative success in estimating compliance independent of shape, they still face limitations in describing compliance across a wide range of materials.
+We aim to create a compliance estimation system that is robust to both shape and material by integrating analytical modeling with machine learning.
+Young’s modulus ($E$) is a fundamental property in mechanical modeling that offers a measure of compliance for nearly all materials [16].
+By parameterizing compliance using Young’s modulus, we can seamlessly incorporate analytical models and create estimates of compliance on a universal scale.
+Our hybrid approach leverages the strengths of analytical modeling to better generalize across materials, while using learning techniques to address complexities posed by diverse geometries.
+Specifically, we use a neural network to fuse CNN-encoded tactile images with estimates from analytical contact models.
+We train our approach across a novel dataset of 285 common objects, including a wide variety of shapes and materials.
+Our primary contribution is a novel compliance estimation archiecture which demonstrates increased accuracy over the current state-of-the-art.
+Additionally, we contribute a novel dataset which could be used in the development of future compliance estimation architectures.
+Using our diverse dataset, we validate that our hybrid estimation approach can generalize well across object shape and material.
+Thus, our estimation system could be applied in general robotic manipulation to characterize unknown objects and inform decision-making.
+This is especially beneficial in scenarios where visually similar objects may have distinct material composition.
 Figure 1: Young’s modulus estimation system.
@@
 Camera-based tactile sensors, such as GelSight sensors, are designed to track contact geometry in high-resolution using a camera placed behind an elastomer gel pad [12, 17, 18, 19, 20, 21].
-They are highly effective in texture recognition [27, 28, 29], shape and pose estimation [30, 31], and measuring contact force [15] and friction [32].
+They are highly effective for sensing texture [22, 23, 24], contact force [12], friction [25], and object shape and pose [26, 27].
 However, reliably modeling contact with these soft sensors remains challenging.
@@
 Finite element analysis has proven a viable way to model soft contact on hardware and used to estimate forces from tactile data [29, 30].
-Though modeling with FEM is accurate, it requires a precise map of object geometry and solutions must be pre-computed offline.
+Though modeling with FEM is accurate, it requires a precise map of object geometry and solutions often must be pre-computed offline.
 Alternative approaches have been developed using Hertzian mechanics to model soft contact in robotic grasping [31, 32].
@@
 Hertzian contact models have been utilized for force estimation [33, 5] and surface curvature detection from tactile data [34].
-Since they are closed-form, these constitutive models can more easily be inverted to determine object compliance.
+These closed-form constitutive models can easily be inverted to determine object compliance.
 2.2 Compliance Estimation
@@
 Similar models have been foundational in developing low-level grasping control policies [35, 36].
-This method can provide an effective stiffness of grasping.
-However, it requires assumptions on object shapes and may struggle when handling complex geometries without a known map of each grasped object.
+This method can provide an effective stiffness of grasping, but may struggle when applied to unknown objects with complex shapes.
 By incorporating tactile sensing, it is possible to create higher fidelity compliance estimation models [37, 38, 39].
 These methods have a stronger ability to sense and generalize across contact geometry but still struggle to account for complex contact geometries.
-Recent learning-based approaches to compliance estimation have demonstrated the ability to generalize across geometry but struggle across a wide-range of material types.
-Work has been done to create a tactile-informed model to estimate the Shore 00 hardness of an object in contact with a GelSight sensor.
-The architecture predicts hardness directly from tactile images without any mechanical modeling [45, 16].
-Shore 00 hardness is a measure designed to quantify the compliance of soft rubbers [46].
-Thus, this method is trained using only rubber objects.
-It is fundamentally restricted to predicting on this scale and cannot express differences in compliance between more rigid objects.
-LLM-backed methods have been developed to classify tactile images with binary semantic descriptions of ‘soft’ or ‘hard’ [12, 13].
-These large-data methods are promising but still have not been used to assess compliance on a continuous scale across material types.
+Recent learning-based approaches to compliance estimation have demonstrated the ability to better generalize across geometry but struggle when applied over a wide-range of material types.
+Data-driven architectures have been developed to estimate the Shore 00 hardness of an object in contact with a GelSight sensor directly from tactile images without any mechanical modeling [40, 14].
+The Shore 00 hardness scale is designed to quantify the compliance of soft rubbers [41].
+This model is trained using exclusively rubber objects and is thus fundamentally limited to this material domain.
+Alternatively, LLM-backed methods have been developed to classify tactile images with binary semantic descriptions of ‘soft’ or ‘hard’ [15, 1].
+These large data models are promising but have not been used to assess finer differences in compliance on a continuous scale.
 Other data-driven implementations have been developed using tactile sensing to directly measure the ripeness of produce [2, 3, 4].
 As before, these methods are restricted by their dataset and labeling method.
-In all cases, labeling choices prevent models from incorporating analytical modeling alongside learning.
+In all cases, labeling choices inhibit models from integrating analytical modeling alongside learning.
 3 Methods
 3.1 Overview
-We develop a hybrid system fusing analytical models alongside learning to estimate the Young’s modulus of general unknown objects from a single parallel grasp.
+We develop a hybrid system that fuses analytical models alongside learning to estimate the Young’s modulus of general unknown objects from a single parallel grasp.
 Through analytical modeling, we can create a well-founded, preliminary estimate for Young’s modulus.
@@
 2.
-Tactile images and measurements of force $F$ and gripper width $w$ are fed into Hertzian and elastic analytical models to fit respective estimates of Young’s modulus $\hat{E}_{\text{elastic}},\hat{E}_{\text{hertz}}$.
-In parallel, a selected number $N_{\text{frames}}$ of tactile images are fed into a data-driven model.
-Features extracted from images are concatenated with grasp measurements into a large fully-connected decoder.
-This multi-tower architecture is similar to other grasp-based learning architectures [49].
-Output features from our learned model are sent to a smaller decoder alongside analytical estimates to produce a final hybrid estimate of Young’s modulus $\hat{E}$.
+After grasping, tactile images and measurements of force $F$ and gripper width $w$ are fed into analytical models to fit respective estimates of Young’s modulus $\hat{E}_{\text{elastic}},\hat{E}_{\text{hertz}}$.
+In parallel, a selected number $N_{\text{frames}}$ of tactile images from a single sensor are processed through a CNN.
+Features extracted from these images are concatenated with grasp measurements $F$, $w$ into a large fully-connected decoder.
+This multi-tower network is similar to other grasp-based architectures [44].
+Output features from the large decoder are sent to a smaller decoder alongside analytical estimates to produce a final hybrid estimate of Young’s modulus $\hat{E}$.
 Figure 2: Young’s modulus estimation architecture.
@@
 3.2 Analytical Contact Models
-Simple Elasticity Parallel grasping can be modeled as uniaxial loading using an adaptation of Hooke’s Law.
+Simple Elasticity: Parallel grasping can be modeled as uniaxial loading using an adaptation of Hooke’s Law.
 A diagram of parallel grasping is provided in Fig.
@@
 Traditional methods have used these measurements alone to estimate the compliance or stiffness of grasped objects [13].
-With camera-based tactile sensing, we can increase the fidelity of the model by measuring contact area $A$ and depth $d$ [15].
+With camera-based tactile sensors, we can increase the fidelity of the model by measuring contact area $A$ and depth $d$ [12].
 Contact area $A$ is determined by masking depth images with constant threshold of 0.1mm.
@@
 Without tactile sensing, we could not measure depth ($d(t)=0$) and would need to assume constant contact area ($A(t)=A_{\text{sensor}}$).
-We use measurements to calculate contact stress $\sigma$ and object strain $\epsilon$, following Eq.
+We calculate contact stress $\sigma$ and object strain $\epsilon$, following Eq.
 1.
+This model assumes symmetry between parallel fingers.
 We define $t=0$ to be the moment of first contact, where $F(t=0)$ must be greater than an empirically-defined threshold.
@@
 $\boldsymbol{\sigma}=\hat{E}_{\text{elastic}}\boldsymbol{\epsilon}$ (2)
-Despite its simplicity, this method does not properly consider the geometry of unknown objects.
-It implicitly assumes objects have a consistent profile along the grasping axis.
-Furthermore, it does not take advantage of precise information about the deformation of contact geometry that is apparent through high-resolution tactile data.
+Despite its simplicity, this method does not properly consider object geometry and implicitly assumes a consistent profile along the grasping axis.
+Furthermore, it overlooks precise information about the deformation of contact geometry that is apparent through high-resolution tactile data.
 Figure 3: Simple mechanics of parallel grasping.
@@
 Timestep $t=0$ is the moment of first contact over the grasp.
-Hertzian Contact Hertzian contact theory examines the deformation of elastic bodies through contact [50].
+Hertzian Contact: Hertzian contact theory models the deformation of elastic bodies in contact [45].
 In a simple case, contact is assumed to be without friction or adhesion [46, 47].
-Unlike the previously-described method, Hertzian mechanics more acutely analyzes geometry through contact.
-It has been used to model soft contact in robotics [36, 39], including GelSight-like tactile sensors [23, 6].
-Precise depth sensing allows us to track surface deformation and apply this model.
+Unlike the previous method, Hertzian mechanics closely analyzes geometry through contact.
+It has been applied to model soft contact in robotics [18, 5], including for GelSight-like tactile sensors [31, 34, 18, 5].
 GelSight sensors are made with highly elastic silicone rubber [17].
-Unless the grasped object is fully rigid, we expect both bodies to deform through contact.
+Unless a grasped object is fully rigid, we expect both bodies to deform through contact.
 This implies the maximum measured penetration depth $d$ will be less than the relative displacement between the finger and object $d<\frac{\Delta w}{2}$.
@@
 3, which depends on the Poisson’s ratio $\nu$ of each body.
-$\frac{1}{E^{*}}=\frac{1-\nu_{\text{sensor}}^{2}}{E_{\text{sensor}}}+\frac{1-\nu_{\text{obj}}^{2}}{E_{\text{obj}}}$ (3)
+$E^{*}=\left(\frac{1-\nu_{\text{sensor}}^{2}}{E_{\text{sensor}}}+\frac{1-\nu_{\text{obj}}^{2}}{E_{\text{obj}}}\right)^{-1}$ (3)
 Hertzian solutions have been derived for a wide range of contact geometries [48].
-Since we are analyzing contact with an unknown object whose shape is not precisely known, we model all contact geometry as spherical.
-This is a reasonable simplification for convex objects, but may produce bounded errors in estimates over our geometrically diverse dataset.
-Our neural network will learn to compensate for this intrinsic and repeatable geometric modeling error.
-Since we assume contact profiles are spherical, we must calculate contact radius $a$.
-This is the radius needed for a circle’s contact area to match the observed contact area $A$.
-Additionally, we can approximate the apparent radius of the unknown object $R$ through observed contact area and gripper displacement [53].
-Using these parameters and the method of dimensionality reduction (MDR), we can calculate the sensor’s surface deformation in response to elastic loading [53, 54].
-This relates force $F$ and depth $d$ by $E^{*}$.
-This relation shows that a harder sphere will plunge deeper into the contact plane, assuming constant force.
-More detailed derivation for this expression and discussion about the MDR is provided in Appendix A.
-$a(t)=\sqrt{\frac{A(t)}{\pi}},\quad R(t)=\frac{a^{2}(t)}{\Delta w(t)},\quad d(t)=\frac{1-\nu_{\text{sensor}}^{2}}{E_{\text{sensor}}}{\left(\frac{3{E^{*}}^{2}F(t)}{32R^{2}(t)}\right)}^{\frac{1}{3}}a(t)$ (4)
-With measured data over the duration of a grasp, we apply least-squares to retrieve a best fit value $\hat{E}^{*}$.
-Then, we use $\hat{E}^{*}$ to compute an estimated modulus for our unknown object $\hat{E}_{\text{hertz}}$, given known properties of our sensor $E_{\text{sensor}}=0.275$ MPa, $\nu_{\text{sensor}}=0.48$.
+We model the sensor as a flat plane.
+Although we are analyzing contact with an object whose shape is unknown, we model the geometry of all grasped objects as spherical.
+This simplification is reasonable for convex objects, but may produce bounded errors over our geometrically diverse dataset.
+Our neural network can learn to compensate for this intrinsic and repeatable geometric modeling error.
+Since we assume contact profiles are spherical, we calculate contact radius $a$ based on observed total contact area $A$.
+Additionally, we calculate the apparent radius of the unknown object $R$ using contact area and gripper displacement [48].
+Using these parameters, we can relate depth $d$ to normal force $F$ via the method of dimensionality reduction (MDR) [48, 49], as formulated in Eq.
+4.
+More detailed derivation for the MDR is provided in Appendix A.
+With these equations established and measurements over the duration of a grasp, we apply least-squares to retrieve a best fit value $\hat{E}^{*}$.
+$a(t)=\sqrt{\frac{A(t)}{\pi}},\quad R(t)=\frac{a^{2}(t)}{\Delta w(t)},\quad\boldsymbol{d}=\frac{1-\nu_{\text{sensor}}^{2}}{E_{\text{sensor}}}{\left(\frac{3{{\hat{E}}^{*2}}\boldsymbol{F}}{32\boldsymbol{R}^{2}}\right)}^{\frac{1}{3}}\boldsymbol{a}$ (4)
+We compute an estimated modulus for our unknown object $\hat{E}_{\text{hertz}}$ through Eq.
+5, given known properties of our sensor $E_{\text{sensor}}=0.275$ MPa, $\nu_{\text{sensor}}=0.48$.
 We assume constant $\nu_{\text{obj}}=0.4$ based on the Poisson’s ratio of common materials [50].
@@
 5 that it will be easiest to discriminate between Young’s moduli nearest to our sensor’s.
-The dataset’s composition is further detailed in Appendix D.
+The dataset’s composition is further detailed in Appendix E.
 As objects were gathered, they were labeled with their Young’s moduli.
-For harder objects, like those made of metals and plastics, this was done by referencing an engineering database of material properties [57], as all objects are made of known materials.
-For softer objects, like those made of rubber, Shore hardness was measured and converted to Young’s modulus.
-Shore hardness is converted to Young’s modulus using Gent’s model of hardness and other established methods [46, 58].
+For harder objects, like those made of plastics and metals, this was done by referencing an engineering database of known material properties [52].
+For softer objects, like those made of rubber, Shore hardness was measured and converted to Young’s modulus using Gent’s model of hardness and other established methods [41, 53].
 These models provide a reliable measure of object compliance on a logarithmic scale.
@@
 5.
-One sensor is inscribed with markers for tracking surface displacement.
-Markers help provide information about shear forces and surface displacement, as further explained in Appendix B.
-By collecting data with different sensors simultaneously, we have the ability to assess the relative necessity of markers in estimating Young’s modulus.
+Although we choose GelSight sensors for our experiment, other flat, camera-based tactile sensors capable of tracking depth in high-resolution would be suitable.
+One sensor is inscribed with markers.
+As explained in Appendix B, markers help enhance the dynamic information in a tactile image by tracking surface displacement.
+Data from only one sensor is required for estimation, but we collect data with both sensors simultaneously to later assess the necessity of markers in estimating Young’s modulus.
+Additionally, we assume symmetry between fingers in our analytical models.
 Normal force is measured through a gauge embedded in the left finger.
@@
 5.2 Data Collection & Training
-Figure 6: Raw force and depth data. Data is collected across grasps for spheres of each material type. Data is fit to functions of the form $d\sim F^{\frac{2}{3}}$, per our Hertzian model.
+Figure 6: Raw force and depth data. Data is collected across grasps for spheres of each material type and fit to $d\sim F^{\frac{2}{3}}$, per our Hertzian model.
 Data Collection: Automated parallel grasps are executed on objects in fixed position.
@@
 Predictions in the gray region are considered sufficiently accurate.
-We train our neural network with preprocessed data as described in Section 5.2.
-Initially, we train a model over only rubber objects.
-This preliminary exercise is performed to evaluate the model’s ability to learn across a constrained dataset.
-Rubber objects are closest to the Young’s modulus of our sensor.
+Initially, we train an instance of our model using only rubber objects from our dataset.
 From Eq.
-5, we expect it will be easier to discern finer differences in compliance between them.
-Results over unseen objects for this model are plotted in Fig.
+5, we expect it will be easier to discern fine differences in compliance between rubber objects, as their Young’s modulus is closest to that of our sensor.
+Estimations across unseen objects for this model are plotted in Fig.
 8.
-The model demonstrates 100% prediction accuracy within an order of magnitude of the ground truth Young’s modulus.
-The prediction range of this experiment is analogous to previously developed Shore hardness tactile models [16], which are designed to estimate the compliance of soft rubber objects.
-Shore 00 hardness values evaluated in [16] range from approximately 6 kPa to 2 MPa.
-Here, we include harder rubber objects as well.
-We train our neural network over the entire collected dataset.
-This learned model is compared to analytical baselines.
-Predictions across the dataset for each method are plotted in Fig.
+The model demonstrates 100% prediction accuracy within an order of magnitude of ground truth Young’s modulus, with an average log10 error of 0.36.
+We utilize this reduced training configuration to compare our architecture against previously developed Shore hardness models [14].
+As detailed in Appendix D, our hybrid system exhibits reduced error over this limited domain.
+Subsequently, we train our neural network over the entire collected dataset and compare against analytical baselines.
+Predictions across unseen objects for each method are plotted in Fig.
 9.
-We observe that as the traditional Hooke’s Law method without tactile sensing fails to distinguish between all material types.
-Simple elastic and Hertzian analytical methods demonstrate significant improvements, but are still less than 30% accurate.
-Our learned model showcases dramatic improvements over all other methods.
-All estimation methods had more difficulty expressing precise differences in Young’s modulus between harder objects that strayed further from the modulus of the sensor (i.e.
-plastic, wood, and metal objects).
-Additionally, we observe decreased overall accuracy from previous evaluation over the rubber subset of our dataset.
-These results are expected given that harder objects are approximately rigid relative to the soft sensor’s Young’s modulus.
-The resolution of the tactile sensor is likely not high enough to detect small differences in deformation between this range of compliances.
+A traditional Hooke’s Law method without tactile sensing fails to distinguish between material types.
+Simple elastic and Hertzian analytical models exhibit stronger performance, but remain less than 30% accurate.
+Our hybrid learned model showcases dramatic improvements over all methods.
 In Table 1, performance is compared between different estimation methods and inputs.
@@
 We evaluate models on a logarithmic scale given the range of Young’s moduli in our dataset.
-Predictions are considered sufficiently accurate within the same order of magnitude of ground truth labels.
 We compute error as the average difference between prediction and ground truth on a log10 scale.
+Predictions are considered accurate within the same order of magnitude of ground truth labels.
+This tolerance is deemed sufficient for downstream tasks of produce sorting based on the change in compliance of selected fruits and vegetables during ripening [54, 55, 56, 57].
 Figure 9: Young’s modulus predictions by material.
@@
 Predictions in the gray region are considered sufficiently accurate.
-From the ablation study, we observe that our learned model vastly reduces error and improves accuracy over analytical methods.
-Furthermore, we see that RGB tactile images with markers are favorable for this estimation task.
-However, we do not observe significant difference in results between models which consider force $F$ and width $w$ measurements.
-This may be due to the fact that all objects were grasped with nearly identical peak force.
-Moreover, force is implicitly detectable through tactile images [16].
-Importantly, considering analytical estimates $\hat{E}_{\text{elastic}},\hat{E}_{\text{hertz}}$ is shown to improve learned prediction results.
-This demonstrates that incorporating analytical modeling improves our ability to discern compliance over a purely learning-based method.
-Accuracy and error metrics are computed across shape and material categories for each method in Appendix D.
+From the ablation study, we observe that our hybrid model drastically out-performs purely analytical and learning-based methods.
+RGB tactile images with markers are shown to be most favorable for this estimation task.
+There is not significant difference in results between models which consider force $F$ and width $w$ measurements.
+This may be due to the fact that all objects were grasped with identical peak force.
+Moreover, force is implicitly apparent through tactile difference images [12].
+Importantly, analytical estimates $\hat{E}_{\text{elastic}},\hat{E}_{\text{hertz}}$ are shown to improve prediction accuracy by nearly 10%.
+This indicates that a hybrid approach to compliance estimation is favorable over purely learning-based methods, which have been previously explored [14, 15, 1, 2, 3, 4].
+Prediction accuracy across shape and material categories for each method is investigated in Appendix E.
 We observe that our model performs independent of shape and can generalize well across materials.
@@
 Table 1: Estimation Architecture Ablation Study
+Limitations: Our estimation system struggled to express precise differences in Young’s modulus between harder objects that stray further from the modulus of our sensor (i.e.
+plastic, wood, and metal objects).
+These results are expected given that harder objects are rigid relative to our soft sensor.
+The resolution of our tactile sensor is not high enough to detect minute differences in deformation between this range of compliances.
+Considering this fact, our estimation system exhibits only marginal effective improvement relative to previously developed Shore hardness estimation models [14].
+Further work must be done to create a truly universal compliance estimation system.
 7 Conclusion
 We have developed a novel system for estimating the Young’s modulus of unknown contacted objects using analytical and data-driven methods.
-This system can create estimates online from only a single parallel grasp.
-It has been shown to perform independent of object shape and demonstrates robustness across materials.
-It can expressively distinguish precise differences in compliance between soft objects.
-The system is less capable of measuring precise differences between harder, such as plastic and metal, objects which are effectively rigid relative to the soft GelSight sensor.
-The estimation system introduced in this paper could be applied for use with other camera-based tactile sensors capable of tracking depth.
+This system can create estimates online from only a single parallel grasp using camera-based tactile sensors.
+It has been shown to perform independent of object shape and remains robust across material types.
+It can precisely distinguish between compliances of soft objects, but struggles to expressively differentiate between that of harder objects.
+The estimation architecture introduced in this paper could be applied for use with other camera-based tactile sensors capable of tracking depth.
 With more diverse tactile data, the accuracy of the system could be further improved.
 Given an increased ability to generalize, this compliance estimation system could be deployed alongside an intelligent high-level controller to inform decision-making in robotic manipulation scenarios by characterizing the compliance of unknown objects.
+This is particularly valuable when visually similar objects may have distinct material composition.
 As one example, it could be utilized to identify the ripeness of fruits or vegetables [1, 2, 4].
@@
 Appendix A Method of Dimensionality Reduction (MDR)
-As discussed in Section 3.2, we model contact between an unknown object and our camera-based tactile sensor as a collision between an elastic sphere and an elastic half-space using Hertzian contact theory.
+As discussed in Section 3.2, we model contact between an unknown object and our flat, camera-based tactile sensor as a collision between an elastic sphere and an elastic half-space using Hertzian contact theory.
 Now, we will walk through the derivation for a constitutive relation using the method of dimensionality reduction (MDR) [58].
@@
 Thus, we can track their relative position to map the displacement of the gel pad surface in both shear and normal directions.
-Markers have been shown to be effective for measuring shear and normal forces from tactile images [15, 22].
-From a broad standpoint, this means we can use markers to increase the amount of dynamic information encoded in a tactile image.
-Without them, we could not estimate shear force.
+Markers have been shown to be effective in estimating shear and normal forces from tactile images [12, 17].
+However, accurate force estimations require assumptions on contact geometry.
+In this work, we seek to estimate the compliance of unknown objects.
+Since we cannot make accurate assumptions on geometry, we measure forces through an external force gauge rather than through our GelSight sensors.
+In any case, we can use markers to increase the amount of dynamic information encoded in a tactile image.
+Without them, we could not perceive shear displacement.
 In our experiment, we have created one GelSight Wedge [17] with markers painted on the surface.
 This was done to investigate how beneficial markers may be for estimating a dynamic property of contact like Young’s modulus.
-The difference between tactile images with or without markers for two objects is tabulated in Fig.
+The difference between tactile images with or without markers for two objects is displayed in Fig.
 11.
@@
 The correlation coefficient between these measurements is computed as 0.84 and bounds are placed at 25% error mark.
-Appendix D Dataset Diversity & Performance
-The diversity of the collected dataset is characterized through categories of material and shape in Tables 3 and 3.
+Appendix D Comparison Against Shore 00 Hardness Baseline
+Previous works have developed learning-based methods to estimate Shore hardness using camera-based tactile sensors without the incorporation of analytical modelling [40, 14, 5, 15, 1].
+Notably, Yuan et al.
+introduced an architecture to estimate Shore 00 hardness of various sample geometries using tactile images captured from a GelSight sensor [14].
+To demonstrate the performance of our novel hybrid estimation system, we directly compare it against results from Yuan et al.
+As discussed in Section 6, we can constrain our dataset to predict compliance for only rubber objects.
+There are 82 rubber objects in our dataset of 285 total objects.
+The compliance of these rubber objects aligns with the Shore 00 hardness scale, which spans approximately from a Young’s modulus of 6 kPa to 2 MPa.
+We can directly evaluate and compare predictions of Young’s modulus with Shore 00 hardness by utilizing established models that convert between the two measures [53].
+Using the rubber subset of our dataset, we replicate experiments conducted by Yuan et al.
+and compare their quoted results against our own [14].
+We use RGB tactile images with markers and analytical estimates as the input to our model.
+First, we train an instance of our hybrid estimation system to predict compliance using only simple shapes such as spheres, cylinders, and rectangular objects.
+This model is validated across seen shapes with unseen compliances.
+Predictions from our model are subsequently converted to Shore 00 hardness to calculate error over the validation set.
+Second, we train another instance of our hybrid model and validate it across unseen objects with arbitrary or irregular shapes.
+It is important to note that Yuan et al.
+evaluate over hand-collected data for the latter experiment, whereas all of our data collection is performed on robot hardware.
+Results from our experiments are presented in Table 2, where root mean-squared error (RMSE) is reported in units of Shore 00 hardness.
+Method Dataset $R^{2}$ RMSE
+Shore 00 Hardness Simple Shapes 0.85 10.3
+Model [14] Arbitrary Shapes 0.39 18.2
+Top-1 Hybrid NN Simple Shapes 0.69 5.9
+(Our Method) Arbitrary Shapes 0.37 14.3
+Table 2: Shore 00 Hardness Prediction Error Comparison
+Our estimation system exhibits decreased error in both experiments across both simple and complex shapes, though the margin of improvement is modest.
+These results suggest that our novel hybrid estimation architecture offers some performance enhancement over previously developed learning-based approaches for predicting the compliance of soft rubbers.
+We can likely attribute this decrease in error to the incorporation of analytical modeling.
+Appendix E Dataset Diversity & Performance
+The diversity of the collected dataset is characterized through categories of material and shape in Tables 4 and 4.
 As intended, the most common material of objects in the dataset is rubber, which includes 28.6% of the objects in the dataset.

### v3

@@
 It can be used to inform low-level contact-rich actions or characterize objects at a high-level.
-In robotic manipulation, existing approaches to estimate compliance have struggled to generalize across object shape and material.
+In robotic manipulation, existing approaches to estimate compliance have struggled to generalize across both object shape and material.
 Using camera-based tactile sensors, proprioception, and force measurements, we present a novel approach to estimate object compliance as Young’s modulus $E$ from parallel grasps.
-We evaluate our method over a novel dataset of 285 common objects, which includes a wide array of shapes and materials with Young’s moduli ranging from 5.0 kPa to 250 GPa.
-Data is collected over automated parallel grasps of each object.
+We evaluate our method over a novel dataset of 285 common objects, including a wide array of shapes and materials with Young’s moduli ranging from 5.0 kPa to 250 GPa.
 Combining analytical and data-driven approaches, we develop a hybrid system using a multi-tower neural network to analyze a sequence of tactile images from grasping.
 This system is shown to estimate the Young’s modulus of unseen objects within an order of magnitude at 74.2% accuracy across our dataset.
-This is a drastic improvement over purely analytical and data-driven baselines which exhibit 28.9% and 65.0% accuracy respectively.
-Importantly, this estimation system performs irrespective of object geometry and demonstrates robustness across object materials.
-Thus, it could be applied in a general robotic manipulation setting to characterize unknown objects and inform decision-making, for instance to sort produce by ripeness.
-Keywords: Robotic Manipulation, Tactile Sensing, Compliance
-1 Introduction
-General and robust object compliance estimation could significantly enhance robotic perception systems.
-For instance, a ripe avocado appears visually identical to an unripe avocado, yet their compliance differs greatly.
-Without the ability to accurately estimate compliance, a robot cannot effectively distinguish between such visually similar objects.
-Compliance measurement has proven valuable in multiple applications, including identifying the ripeness of fruits and vegetables [1, 2, 3, 4] and in biomedical fields [5, 6, 7].
-In the future, a universal compliance estimation architecture could be applied in assembly or recycling processes to inform decisions based on object material.
-In robotics, we seek to replicate the human ability to perceive properties like compliance through touch [8].
-Humans naturally use tactile information to characterize objects and inform interactions [9].
-In the case of compliance, we sense discrepancies between soft and hard objects through fingertip pressure and deformation [10].
-High-resolution, camera-based tactile sensors like GelSight are specifically designed to track the deformation of a robotic fingertip during contact [11, 12].
-However, even with the help of high-resolution tactile sensing, achieving human-like compliance perception remains challenging, especially in unstructured environments.
-An object’s deformation in response to loading depends on both its shape and material composition.
-Thus, without known geometry and pose tracking, analytical methods cannot be utilized to rigorously determine compliance [13].
-Alternatively, learning-based implementations have been developed to estimate compliance from tactile images alone without the use of mechanical modeling [14, 5, 15, 1].
-While these methods have demonstrated relative success in estimating compliance independent of shape, they still face limitations in describing compliance across a wide range of materials.
-We aim to create a compliance estimation system that is robust to both shape and material by integrating analytical modeling with machine learning.
-Young’s modulus ($E$) is a fundamental property in mechanical modeling that offers a measure of compliance for nearly all materials [16].
-By parameterizing compliance using Young’s modulus, we can seamlessly incorporate analytical models and create estimates of compliance on a universal scale.
-Our hybrid approach leverages the strengths of analytical modeling to better generalize across materials, while using learning techniques to address complexities posed by diverse geometries.
-Specifically, we use a neural network to fuse CNN-encoded tactile images with estimates from analytical contact models.
-We train our approach across a novel dataset of 285 common objects, including a wide variety of shapes and materials.
-Our primary contribution is a novel compliance estimation archiecture which demonstrates increased accuracy over the current state-of-the-art.
-Additionally, we contribute a novel dataset which could be used in the development of future compliance estimation architectures.
-Using our diverse dataset, we validate that our hybrid estimation approach can generalize well across object shape and material.
-Thus, our estimation system could be applied in general robotic manipulation to characterize unknown objects and inform decision-making.
-This is especially beneficial in scenarios where visually similar objects may have distinct material composition.
+This is an improvement over purely analytical and data-driven baselines which exhibit 28.9% and 65.0% accuracy respectively.
+Importantly, this estimation system performs irrespective of object geometry and demonstrates increased robustness across material types.
+I Introduction
+Humans naturally use tactile feeedback to characterize objects and inform our interactions with them [1].
+Compliance, or softness, is useful in everyday tasks.
+For example, we can determine the ripeness of an avocado or other fruits by assessing their compliance.
+Crucially, surgeons rely on tactile feedback and perceived compliance to accurately characterize tissues and detect tumors [2].
+In robotics, if we would like to better replicate the natural proficiency of human manipulation through teleoperated action, we must communicate this tactile information to remote users.
+Recently, haptic devices have been developed to display softness for remote users [3, 4, 5].
+However, these displays are only useful if we can reliably provide them a measure of compliance in real-time.
 Figure 1: Young’s modulus estimation system.
@@
 Tactile images are size 35mm by 25mm.
-2 Related Work
-2.1 GelSight Sensors & Contact Modeling
-Camera-based tactile sensors, such as GelSight sensors, are designed to track contact geometry in high-resolution using a camera placed behind an elastomer gel pad [12, 17, 18, 19, 20, 21].
-They are highly effective for sensing texture [22, 23, 24], contact force [12], friction [25], and object shape and pose [26, 27].
+Humans sense discrepancies between soft and hard objects through fingertip pressure and deformation [6].
+Camera-based tactile sensors, like GelSight, are specifically designed to track these parameters on robotic fingertips [7, 8].
+Analytical models, such as Hertzian mechanics, follow a similar approach to approximate compliance as humans may [9, 10].
+Ideally, we could input tactile data to these models to determine compliance.
+However, traditional analytical frameworks often struggle to generalize across arbitrary geometries that may be encountered in unstructured environments.
+Since an object’s deformation under load depends on both its shape and compliance, analytical methods cannot rigorously determine compliance without prior knowledge of object geometry [11].
+Alternatively, learning-based implementations have been developed to estimate compliance from tactile images alone without the use of mechanical modeling [12, 13, 14, 15].
+While these methods have demonstrated success in estimating compliance independent of shape, they still face limitations in describing compliance across a wide range of materials.
+Many of these models elect to exclusively label the compliance of rubber objects via the Shore hardness scale.
+To successfully incorporate compliance feedback into a teleoperation interface, we need labels of compliance on a universal scale to replicate the continuity that humans experience in real-world manipulation.
+To address these limitations and move towards more generalized compliance estimation, we propose a hybrid approach that combines analytical modeling with learning-based methods to estimate an object’s compliance as Young’s modulus $E$ from a single parallel grasp.
+By incorporating analytical models, we improve generalization across materials, while using a learned CNN to account for geometric complexities.
+We choose to label compliance as Young’s modulus to align with analytical models and so that our system produces estimations on a universal scale.
+In this work, we present the following contributions:
+  • A novel compliance estimation architecture which demonstrates increased accuracy over the current state-of-the-art.
+Unlike previous methods, our hybrid learned architecture incorporates analytical modeling.
+  • Introduce Young’s modulus $E$ as a label of compliance to enable seamless incorporation of analytical models and produce estimates on a universal scale.
+  • A novel dataset containing 285 common objects, featuring a wide variety of shapes and materials, which could be used in developing future compliance estimation architectures.
+We validate that our hybrid estimation approach can generalize reasonably well across our diverse dataset.
+Thus, our estimation system could be applied in robotic manipulation systems to more robustly estimate compliance.
+II Related Works
+II-A GelSight Sensors & Contact Modeling
+Camera-based tactile sensors, such as GelSight, are designed to track contact geometry in high-resolution using a camera placed behind an elastomer gel pad [8, 16, 17].
+They are effective for sensing texture [18], object shape and pose [19], contact force [8], and friction [20].
 However, reliably modeling contact with these soft sensors remains challenging.
-Simulators have been developed to generate tactile images and deformed geometry of sensors using pre-computed finite element method (FEM) solutions [28].
-Finite element analysis has proven a viable way to model soft contact on hardware and used to estimate forces from tactile data [29, 30].
-Though modeling with FEM is accurate, it requires a precise map of object geometry and solutions often must be pre-computed offline.
-Alternative approaches have been developed using Hertzian mechanics to model soft contact in robotic grasping [31, 32].
-This closed-form method offers improved computational efficiency over FEM and simulated methods.
+Finite element method (FEM) has proven a viable option [21, 22].
+Although modeling with FEM is accurate, it requires a precise map of object geometry and solutions often must be pre-computed offline.
+Alternative approaches have been developed using Hertzian mechanics to model soft contact in robotic grasping [23].
+This closed-form method offers improved computational efficiency over FEM and other methods.
 Still, it requires assumptions on contact geometry, namely that geometry should be axisymmetric.
-Hertzian contact models have been utilized for force estimation [33, 5] and surface curvature detection from tactile data [34].
-These closed-form constitutive models can easily be inverted to determine object compliance.
-2.2 Compliance Estimation
-Estimating the compliance of a contacted object remains a complex problem in robotic manipulation, particularly without assumptions on object shape or material type.
+Hertzian contact models have been utilized for force estimation [24, 13] and curvature detection from tactile data [25].
+These closed-form constitutive models can be inverted to determine compliance.
+II-B Compliance Estimation
+Estimating the compliance of a contacted object remains a difficult problem in robotic manipulation, particularly without assumptions on object shape or material type.
 Analytical approaches to compliance estimation often lack robustness to variations in geometry.
-A traditional approach to model compliance is to apply Hooke’s Law using measurements of grasping force and positions [13].
-Similar models have been foundational in developing low-level grasping control policies [35, 36].
+The most traditional approach to model compliance is to apply Hooke’s Law to measurements of grasping force and position [11].
 This method can provide an effective stiffness of grasping, but may struggle when applied to unknown objects with complex shapes.
-By incorporating tactile sensing, it is possible to create higher fidelity compliance estimation models [37, 38, 39].
-These methods have a stronger ability to sense and generalize across contact geometry but still struggle to account for complex contact geometries.
-Recent learning-based approaches to compliance estimation have demonstrated the ability to better generalize across geometry but struggle when applied over a wide-range of material types.
-Data-driven architectures have been developed to estimate the Shore 00 hardness of an object in contact with a GelSight sensor directly from tactile images without any mechanical modeling [40, 14].
+By incorporating tactile sensing, it is possible to create higher fidelity compliance estimation models [26, 27].
+These methods have a stronger ability to generalize across contact geometry through sensing, but still struggle to account for complex contact geometries.
+Recent learning-based approaches to compliance estimation have demonstrated the ability to better generalize across geometry, but struggle when applied over a wide-range of material types.
+Data-driven architectures have been developed to estimate the Shore 00 hardness of an object in contact with a GelSight sensor directly from tactile images without any mechanical modeling [12].
 The Shore 00 hardness scale is designed to quantify the compliance of soft rubbers [28].
-This model is trained using exclusively rubber objects and is thus fundamentally limited to this material domain.
-Alternatively, LLM-backed methods have been developed to classify tactile images with binary semantic descriptions of ‘soft’ or ‘hard’ [15, 1].
-These large data models are promising but have not been used to assess finer differences in compliance on a continuous scale.
-Other data-driven implementations have been developed using tactile sensing to directly measure the ripeness of produce [2, 3, 4].
+This model is trained using exclusively rubber objects and is thus limited to this material domain.
+Alternatively, LLM-backed methods have been developed to classify tactile images with binary semantic descriptions of “soft” or “hard” [14, 15].
+These large data models are promising, but have not been used to assess finer differences in compliance on a continuous or quantitative scale.
+Other data-driven implementations have been developed using tactile sensing to directly detect the ripeness of produce [29, 30, 31].
 As before, these methods are restricted by their dataset and labeling method.
-In all cases, labeling choices inhibit models from integrating analytical modeling alongside learning.
-3 Methods
-3.1 Overview
+In all cases, models do not incorporate analytical modeling alongside learning.
+Figure 2: Young’s modulus estimation architecture.
+Tactile images, depth, and measured grasping force $F$ and width $w$ are sent into analytical and learned models.
+Estimates from analytical models are fed into a fully-connected decoder alongside learned grasp features to produce hybrid estimate $\hat{E}$.
+III Methods
+III-A Overview
 We develop a hybrid system that fuses analytical models alongside learning to estimate the Young’s modulus of general unknown objects from a single parallel grasp.
@@
 A large, diverse dataset of 285 common objects is collected to train and evaluate our method.
-Our approach of combining analytical models with data-driven methods is partially inspired by residual physics architectures [42] and physics-informed tactile models [43].
-Our hybrid estimation architecture is displayed in Fig.
+This approach of combining analytical models with data-driven methods is partially inspired by residual physics architectures [32] and physics-informed tactile models [33].
+Our hybrid estimation architecture is shown in Fig.
 2.
-After grasping, tactile images and measurements of force $F$ and gripper width $w$ are fed into analytical models to fit respective estimates of Young’s modulus $\hat{E}_{\text{elastic}},\hat{E}_{\text{hertz}}$.
-In parallel, a selected number $N_{\text{frames}}$ of tactile images from a single sensor are processed through a CNN.
-Features extracted from these images are concatenated with grasp measurements $F$, $w$ into a large fully-connected decoder.
-This multi-tower network is similar to other grasp-based architectures [44].
-Output features from the large decoder are sent to a smaller decoder alongside analytical estimates to produce a final hybrid estimate of Young’s modulus $\hat{E}$.
-Figure 2: Young’s modulus estimation architecture.
-Tactile images, depth, and measured grasping force $F$ and width $w$ are sent into analytical and learned models.
-Estimates from analytical models are fed into a fully-connected decoder alongside learned grasp features.
-Code for the project is publicly available on GitHub. Collected data for multiple grasps of all objects with labeled Young’s moduli and Shore hardness is publicly available on HuggingFace.
-3.2 Analytical Contact Models
-Simple Elasticity: Parallel grasping can be modeled as uniaxial loading using an adaptation of Hooke’s Law.
+After grasping, tactile images and measurements of normal force $F$ and gripper width $w$ are input into analytical models to fit respective estimates of Young’s modulus $\hat{E}_{\text{elastic}},\hat{E}_{\text{hertz}}$.
+In parallel, a selected number $N_{\text{frames}}$ of tactile images from a single sensor are passed to a CNN.
+Our CNN architecture is adapted from [34], which has demonstrated strong performance in classifying sequences of tactile images.
+Features extracted from tactile images by this CNN are concatenated with grasp measurements $F$, $w$ and fed into a large fully-connected decoder.
+Output features from the large decoder are sent alongside analytical estimates to a smaller decoder to produce a final hybrid estimate of Young’s modulus $\hat{E}$.
+This multi-tower network is inspired by other grasp-based architectures [35].
+Code for the project is publicly available on GitHub. Collected data with labeled Young’s moduli and Shore hardnesses is publicly available on HuggingFace.
+III-B Analytical Contact Models
+Simple Elasticity: Parallel grasping can be modeled as uniaxial loading following Hooke’s Law.
 A diagram of parallel grasping is provided in Fig.
+3a.
+At every instance $t$, normal force $F$ is applied to the grasped object, with gripper at width $w$.
+Traditional methods have used these measurements alone to estimate the stiffness of grasped objects [11].
+Throughout this paper, we assume objects to be made of isotropic materials.
+With camera-based tactile sensors, we can increase the fidelity of contact models by measuring contact area $A$ and depth $d$ [8].
+Contact area $A$ is determined by masking tactile depth images with a constant threshold of 0.1mm.
+This threshold was chosen empirically to cancel out potential noise.
+Without tactile sensing, we could not measure depth ($d(t)=0$) and would need to assume a constant contact area ($A(t)=A_{\text{sensor}}$).
+We calculate contact stress $\sigma$ and object strain $\epsilon$, following Eq.
+1 and Eq.
+2.
+This model assumes symmetry between parallel fingers.
+We define $t=0$ to be the moment of first contact, where $F(t=0)$ must be greater than 0.75 N.
+$\sigma(t)=\frac{F(t)}{A(t)}$ (1)
+$\epsilon(t)=\frac{\Delta w(t)+2\Delta d(t)}{w(0)+2d(0)}=\frac{(w(t)+2d(t))-(w(0)+2d(0))}{w(0)+2d(0)}$ (2)
+Young’s modulus is defined as the ratio between elastic stress and strain [36].
+Using measurements across grasping time $t$, we can fit a linear estimate for Young’s modulus $\hat{E}_{\text{elastic}}$ following Eq.
 3.
-At every instance $t$, force $F$ is applied to the grasped object, with gripper at width $w$.
-Traditional methods have used these measurements alone to estimate the compliance or stiffness of grasped objects [13].
-With camera-based tactile sensors, we can increase the fidelity of the model by measuring contact area $A$ and depth $d$ [12].
-Contact area $A$ is determined by masking depth images with constant threshold of 0.1mm.
-This threshold was chosen empirically to cancel out potential noise.
-Without tactile sensing, we could not measure depth ($d(t)=0$) and would need to assume constant contact area ($A(t)=A_{\text{sensor}}$).
-We calculate contact stress $\sigma$ and object strain $\epsilon$, following Eq.
-1.
-This model assumes symmetry between parallel fingers.
-We define $t=0$ to be the moment of first contact, where $F(t=0)$ must be greater than an empirically-defined threshold.
-$\sigma(t)=\frac{F(t)}{A(t)}\;,\quad\quad\epsilon(t)=\frac{\Delta w(t)+2\Delta d(t)}{w(0)+2d(0)}=\frac{(w(t)+2d(t))-(w(0)+2d(0))}{w(0)+2d(0)}$ (1)
-In the elastic regime, stress $\sigma$ and $\epsilon$ are related by Young’s modulus [16].
-Using measurements across grasping time $t$, we fit an estimate for Young’s modulus $\hat{E}_{\text{elastic}}$ following Eq.
-2.
-$\boldsymbol{\sigma}=\hat{E}_{\text{elastic}}\boldsymbol{\epsilon}$ (2)
-Despite its simplicity, this method does not properly consider object geometry and implicitly assumes a consistent profile along the grasping axis.
+$\bm{\sigma}=\hat{E}_{\text{elastic}}\bm{\epsilon}$ (3)
+Despite its simplicity, this method does not accurately calculate object stress or strain, especially if the object is larger than the sensor surface.
 Furthermore, it overlooks precise information about the deformation of contact geometry that is apparent through high-resolution tactile data.
-Figure 3: Simple mechanics of parallel grasping.
-Spherical object of radius $R$ is grasped with soft GelSight sensors.
+Figure 3: Parallel grasping and Hertzian mechanics.
+(a) Sphere of radius $R$ is grasped with soft GelSight sensors.
 Normal force $F$ is applied at width $w$.
-Penetration depth $d$ and contact radius $a$ can be measured with tactile sensors.
+Tactile penetration depth $d$ and contact radius $a$ are measured with tactile sensors.
 Timestep $t=0$ is the moment of first contact over the grasp.
-Hertzian Contact: Hertzian contact theory models the deformation of elastic bodies in contact [45].
+(b) Normal force $F$ and maximum tactile depth $d$ measurements are collected across grasps for spheres of each material type and fit to $d\sim F^{\frac{2}{3}}$, per our Hertzian model.
+Hertzian Contact: Hertzian contact theory models the deformation of elastic bodies in contact with each other.
 In a simple case, contact is assumed to be without friction or adhesion [10, 9].
-Unlike the previous method, Hertzian mechanics closely analyzes geometry through contact.
-It has been applied to model soft contact in robotics [18, 5], including for GelSight-like tactile sensors [31, 34, 18, 5].
-GelSight sensors are made with highly elastic silicone rubber [17].
-Unless a grasped object is fully rigid, we expect both bodies to deform through contact.
-This implies the maximum measured penetration depth $d$ will be less than the relative displacement between the finger and object $d<\frac{\Delta w}{2}$.
-This phenomenon is depicted in Fig.
-3.
+Unlike previous methods, Hertzian mechanics closely analyzes geometry through contact.
+It has poreviously been applied to model soft contact in robotics with camera-based tactile sensors [23, 25, 22, 13].
+GelSight sensors are made with elastic silicone rubber [16].
+Unless a grasped object is rigid, we expect both bodies to deform through contact.
+This implies the maximum measured tactile depth $d$ will be less than the relative displacement between the finger and object $d<\frac{\Delta w}{2}$, as depicted in Fig.
+3a.
 An aggregate modulus $E^{*}$ describes the bodies’ combined resistance to deformation in Eq.
-3, which depends on the Poisson’s ratio $\nu$ of each body.
+4, which depends on the Poisson’s ratio $\nu$ of each body.
 $E^{*}=\left(\frac{1-\nu_{\text{sensor}}^{2}}{E_{\text{sensor}}}+\frac{1-\nu_{\text{obj}}^{2}}{E_{\text{obj}}}\right)^{-1}$ (4)
 Hertzian solutions have been derived for a wide range of contact geometries [37].
-We model the sensor as a flat plane.
-Although we are analyzing contact with an object whose shape is unknown, we model the geometry of all grasped objects as spherical.
-This simplification is reasonable for convex objects, but may produce bounded errors over our geometrically diverse dataset.
-Our neural network can learn to compensate for this intrinsic and repeatable geometric modeling error.
+We modeled the sensor as a half-space, which may be a reasonable approximation in most cases but could be less accurate for large, rigid objects that penetrate deeper into the sensor’s finite thickness.
+We modeled all grasped objects as spherical.
+This assumption is reasonable for convex objects, but may produce bounded errors over our geometrically diverse dataset.
+We expect that out neural network can learn to compensate for this intrinsic and repeatable geometric modeling error.
 Since we assume contact profiles are spherical, we calculate contact radius $a$ based on observed total contact area $A$.
 Additionally, we calculate the apparent radius of the unknown object $R$ using contact area and gripper displacement [37].
-Using these parameters, we can relate depth $d$ to normal force $F$ via the method of dimensionality reduction (MDR) [48, 49], as formulated in Eq.
+Using these parameters, we can relate tactile depth $d$ to normal force $F$ as formulated in Eq.
+6.
+This relationship is derived via the method of dimensionality reduction (MDR) in Appendix -A [9, 37, 38].
+With equations established and measurements over the duration of a grasp, we can apply least-squares to retrieve a best fit value $\hat{E}^{*}$.
+$a(t)=\sqrt{\frac{A(t)}{\pi}},\quad R(t)=\frac{a^{2}(t)}{\Delta w(t)}$ (5)
+$\bm{d}=\frac{1-\nu_{\text{sensor}}^{2}}{E_{\text{sensor}}}{\left(\frac{3{{\hat{E}}^{*2}}\bm{F}}{32\bm{R}^{2}}\right)}^{\frac{1}{3}}\bm{a}$ (6)
+With calculated $\hat{E}^{*}$, we can extract an estimated modulus for an unknown object $\hat{E}_{\text{hertz}}$ through Eq.
+7, given known properties of our sensor $E_{\text{sensor}}=0.275$ MPa, $\nu_{\text{sensor}}=0.48$.
+We assume constant $\nu_{\text{obj}}=0.4$ based on the Poisson’s ratio of common materials [39].
+After estimates are computed, a linear scaling is added to account for potential constant errors across the dataset.
+$\hat{E}_{\text{hertz}}=\left(1-\nu_{\text{obj}}^{2}\right){\left(\frac{1}{\hat{E}^{*}}-\frac{1-\nu_{\text{sensor}}^{2}}{E_{\text{sensor}}}\right)}^{-1}$ (7)
+IV Physical Dataset
+A novel physical dataset of 285 common household objects was gathered to train and validate methods.
+For comparison, this is nearly 4 times larger than the YCB dataset, which includes only 77 objects [40].
+Additionally, each object was required to have approximately uniform material composition to ensure that an accurate Young’s modulus could be defined for the object and used to label data for supervised training.
+Furthermore, chosen objects for training must not exhibit viscoelastic behavior.
+A subset of total dataset objects is displayed in Fig.
 4.
-More detailed derivation for the MDR is provided in Appendix A.
-With these equations established and measurements over the duration of a grasp, we apply least-squares to retrieve a best fit value $\hat{E}^{*}$.
-$a(t)=\sqrt{\frac{A(t)}{\pi}},\quad R(t)=\frac{a^{2}(t)}{\Delta w(t)},\quad\boldsymbol{d}=\frac{1-\nu_{\text{sensor}}^{2}}{E_{\text{sensor}}}{\left(\frac{3{{\hat{E}}^{*2}}\boldsymbol{F}}{32\boldsymbol{R}^{2}}\right)}^{\frac{1}{3}}\boldsymbol{a}$ (4)
-We compute an estimated modulus for our unknown object $\hat{E}_{\text{hertz}}$ through Eq.
-5, given known properties of our sensor $E_{\text{sensor}}=0.275$ MPa, $\nu_{\text{sensor}}=0.48$.
-We assume constant $\nu_{\text{obj}}=0.4$ based on the Poisson’s ratio of common materials [50].
-After analytical estimates are computed, linear scaling is added to account for potential constant errors across the dataset.
-$\hat{E}_{\text{hertz}}=\left(1-\nu_{\text{obj}}^{2}\right){\left(\frac{1}{\hat{E}^{*}}-\frac{1-\nu_{\text{sensor}}^{2}}{E_{\text{sensor}}}\right)}^{-1}$ (5)
-4 Physical Dataset
-A novel physical dataset of 285 objects was gathered to train and validate our method.
-For comparison, our dataset is nearly 4 times larger than the YCB dataset, which includes only 77 objects [51].
-Chosen objects were intended to be among things a robotic manipulator may commonly interact with in general operation.
-Additionally, each object was required to have approximately uniform material composition to ensure that an accurate Young’s modulus could be defined for the object and used to label data for supervised training.
-A subset of dataset objects is displayed in Fig.
-4.
-Figure 4: Physical dataset.
-(a) Selected objects from our physical dataset are displayed.
-These are common objects with a wide variety of shapes and materials.
-(b) Some custom objects are molded out of silicone to replicate the shape of non-rubber objects in the dataset.
-To evaluate how well estimation algorithms could generalize, diversity of shape and material was prioritized in acquiring objects.
-Thus, we mold 20 custom objects out of silicone to replicate the exact shape of other objects in the dataset with vastly different material composition.
+Some objects, such as bananas and other real fruit, are not used for model training.
+Figure 4: Physical dataset. (a) Selected objects from our physical dataset are displayed. (b) Custom objects are molded out of silicone to replicate the shape of non-rubber objects in the dataset.
+To evaluate estimation algorithms’ ability to generalize, diversity of shape and material were prioritized in acquiring objects.
+We mold 20 custom objects out of silicone to replicate the exact shape of objects in the dataset that are made of much harder materials.
 Molded objects are displayed in Fig.
 4b.
-Training on data collected across object replicas of different materials will help prevent overfitting to shapes and textures.
-Moreover, there is incentive to bias towards rubber objects because it is clear from Eq.
-5 that it will be easiest to discriminate between Young’s moduli nearest to our sensor’s.
-The dataset’s composition is further detailed in Appendix E.
-As objects were gathered, they were labeled with their Young’s moduli.
-For harder objects, like those made of plastics and metals, this was done by referencing an engineering database of known material properties [52].
-For softer objects, like those made of rubber, Shore hardness was measured and converted to Young’s modulus using Gent’s model of hardness and other established methods [41, 53].
-These models provide a reliable measure of object compliance on a logarithmic scale.
-The dataset includes materials ranging from foam to steel, with Young’s moduli from 5.0 kPa to 250 GPa.
+Training on data collected across object replicas of different materials may help prevent overfitting to shapes and textures.
+The dataset’s composition is detailed in Tables II and II.
+TABLE I: Dataset Materials
+Material Percent (%)
+Rubber 28.6%
+Metal 25.1%
+Plastic 24.0%
+Wood 8.4%
+Foam 6.3%
+Other 7.6%
+TABLE II: Dataset Shapes
+Shape Percent (%)
+Cylinder 20.2%
+Sphere 11.5%
+Rectangular 15.3%
+Hexagonal 2.1%
+Irregular 50.9%
+As objects were gathered, they were labeled with their respective Young’s moduli.
+Hard objects, made of plastics, wood, glass, and metals, were labeled by referencing an engineering database of known material properties [41].
+This was possible given that all selected hard objects are made of known materials.
+While variations due to manufacturing conditions or other factors are possible, these labels are expected to be relatively accurate with respect to our vast material domain.
+Soft objects, made of rubber and foam, were labeled by converting Shore hardness measurements to Young’s modulus using Gent’s model and other conversion methods [28, 42, 43].
+Each soft object was indented with a durometer 10 times and the median measurement was taken.
+Conversion models provide a reliable measure of object compliance on a logarithmic scale, making them a widely used approximation in material characterization [3].
+To further validate assigned mechanical properties, 7 rubber and foam objects, including our sensor’s gel pad, were probed with a Dynamic Mechanical Analyzer (TA Instrument Q850) over a force range of 0 to 10 N at a rate of 2 N/min.
+Results from this testing are plotted in Fig.
+5 confirming the accuracy of our labeling techniques.
+In total, the dataset includes materials ranging from foam to steel, with Young’s moduli from 5.0 kPa to 250 GPa.
 For reference, our sensor’s Young’s modulus is 0.275 MPa.
-5 Experiments
-5.1 Hardware Setup
-Figure 5: Hardware setup for data collection. GelSight Wedge sensors are attached to Franka Panda gripper with embedded force gauge.
-GelSight Wedge sensors [17] were rigidly attached to a standard parallel robotic gripper with 3D-printed plastic fingers, as depicted in Fig.
-5.
-Although we choose GelSight sensors for our experiment, other flat, camera-based tactile sensors capable of tracking depth in high-resolution would be suitable.
-One sensor is inscribed with markers.
-As explained in Appendix B, markers help enhance the dynamic information in a tactile image by tracking surface displacement.
-Data from only one sensor is required for estimation, but we collect data with both sensors simultaneously to later assess the necessity of markers in estimating Young’s modulus.
+Figure 5: Dynamic Mechanical Analyzer results.
+Compressive stress-strain curves recorded on samples of selected foam and rubber objects from our dataset.
+Dotted lines respresent expected linear curves from labeled Young’s modulus $E$ for each object, where moduli were approximated via Shore hardness.
+Figure 6: Hardware for data collection and example tactile data.
+(a) GelSight Wedge sensors are attached to Franka Panda gripper with embedded force gauge.
+(b) Tactile images and depth reconstruction from a GelSight Wedge sensor without markers over a set of example objects.
+Shown frames are all from peak pressing force of 30 N.
+All displayed tactile images are size 35mm by 25mm.
+V Data Collection & Training
+V-A Hardware Setup
+GelSight Wedge sensors [16] were rigidly attached to a Franka Panda parallel robotic gripper with 3D-printed fingers, as depicted in Fig.
+6a.
+Fingers were printed using FormLabs Tough 2000 resin.
+Although we used GelSight sensors for our experiment, other camera-based tactile sensors capable of tracking contact depth in high-resolution would be suitable.
+One sensor was inscribed with black dot markers to help enhance the dynamic information in a tactile image by tracking surface displacement [8, 16].
+Data from only one sensor is required for estimation, but we collected data from both sensors simultaneously to later assess the necessity of markers in estimating Young’s modulus.
 Additionally, we assume symmetry between fingers in our analytical models.
-Normal force is measured through a gauge embedded in the left finger.
-Gripper position is directly measured from control software.
-5.2 Data Collection & Training
-Figure 6: Raw force and depth data. Data is collected across grasps for spheres of each material type and fit to $d\sim F^{\frac{2}{3}}$, per our Hertzian model.
-Data Collection: Automated parallel grasps are executed on objects in fixed position.
-For each object in our dataset, a set number of grasps are recorded.
-Objects are re-oriented between each grasp to diversify potential antipodal grasp locations.
-During grasping, tactile images from GelSight sensors, forces, and gripper widths, are recorded at 30 Hz.
-Measurements are synchronized by shifting for latency between cameras.
-Grasps are executed at constant velocity up to force of 30 N.
+Normal force $F$ was measured through an HX711 force gauge in the left finger.
+Gripper position was obtained directly from control software.
+To improve accuracy, measured force was used to correct gripper position readings by compensating for finger deflection with a fitted linear model.
+V-B Data Collection & Training
+Data Collection: Automated parallel grasps were executed on objects in fixed position.
+For each object in our dataset, a set number of grasps were recorded.
+Objects were re-oriented between each grasp to diversify potential antipodal grasp locations.
+During grasping, tactile images from GelSight sensors, normal forces, and gripper widths, were recorded at 30 Hz.
+Measurements were synchronized by shifting for latency between cameras.
+Grasping was executed at constant velocity of 3.75 cm/s and held at closed position for 0.5 s.
+Peak normal force varied across each grasp, up to 30 N.
 Raw data collected from grasps of spherical objects is plotted in Fig.
-6.
-More details about processing data collected from grasps is provided in Appendix C.
+3b.
 Tactile and depth images at grasping force of 30 N for a set of example objects are depicted in Fig.
+6b.
+As expected from Eq.
+6, peak depth is higher for objects of the same shape with higher modulus.
+Additionally, textures are more finely captured for rigid objects with higher modulus.
+Data Preprocessing: Data recorded from grasping objects was clipped to include only the loading sequence, spanning from initial contact to peak force.
+Misaligned grasps where loading is non-monotonic were disregarded.
+A fixed number of frames ($N_{\text{frames}}=3$) were sampled equidistantly in time across the loading sequence.
+To enhance training diversity, multiple sets of these samples were created from each grasp by shifting the starting point by one timestep.
+As illustrated in Fig.
+2, the sampled tactile images, along with analytical estimates, normal force $F$, and width $w$ measurements, were sent as inputs to our multi-tower architecture.
+Additionally, the images undergo random horizontal and vertical flipping for data augmentation.
+Training: We trained our model for 80 episodes over 4,000 input grasps.
+In total, these grasps include over 12,000 tactile images.
+Training was performed end-to-end without any pretraining.
+Output labels were normalized on a log10 scale given the range of Young’s moduli in our dataset.
+Final sigmoid activation function was used with an MSE loss function.
+VI Results
+Initially, we train an instance of our model using only rubber objects from our dataset.
+From Eq.
+7, it is expected to be easier to discern fine differences in compliance between rubber objects, as their Young’s modulus is closest to that of our sensor.
+Estimations across unseen objects for this model are plotted in Fig.
 7.
-We observe that textures are more finely captured for rigid object’s with higher modulus.
-As expected by Eq.
-4, peak depth is higher for objects of the same shape with higher modulus.
-Data Preprocessing: Data recorded from grasping objects is clipped to include only the loading sequence from first contact to peak force.
-Misaligned grasps where loading is non-monotonic are disregarded.
-From here, a set number of frames $N_{\text{frames}}=3$ are sampled equidistantly in time.
-Multiple augmentations of samples are taken from each grasp.
-As depicted in Fig.
-2, tactile images are sent as inputs to our multi-tower architecture alongside analytical estimates and force and width measurements.
-Images are randomly transformed with horizontal and vertical flipping.
-Figure 7: Example tactile data.
-Tactile images and depth reconstruction from a GelSight Wedge sensor without markers over a set of example objects.
-Shown frames are all from peak pressing force of 30 N.
-Tactile images are size 35mm by 25mm.
-Training: We train our model for 80 iterations over 4,000 grasp inputs.
-In total, these grasps include more than 12,000 tactile images.
-Output labels are normalized on a log10 scale given the range of Young’s moduli in our dataset.
-Final sigmoid activation function is used with an MSE loss function.
-6 Experimental Results
+The model demonstrates 100% prediction accuracy within an order of magnitude of ground truth Young’s modulus, with an average log10 error of 0.36.
 Figure 7: Young’s modulus predictions for rubber objects only.
 Results for model trained on only rubber objects over 80 randomly sampled grasps of unseen objects.
+Predictions in the gray region are considered sufficiently accurate within an order of magnitude of ground truth.
+Figure 8: Young’s modulus predictions and model performance.
+(a) Results for learned and analytical estimation methods are plotted, using predictions from 80 randomly sampled grasps per material category.
 Predictions in the gray region are considered sufficiently accurate.
-Initially, we train an instance of our model using only rubber objects from our dataset.
-From Eq.
-5, we expect it will be easier to discern fine differences in compliance between rubber objects, as their Young’s modulus is closest to that of our sensor.
-Estimations across unseen objects for this model are plotted in Fig.
-8.
-The model demonstrates 100% prediction accuracy within an order of magnitude of ground truth Young’s modulus, with an average log10 error of 0.36.
-We utilize this reduced training configuration to compare our architecture against previously developed Shore hardness models [14].
-As detailed in Appendix D, our hybrid system exhibits reduced error over this limited domain.
-Subsequently, we train our neural network over the entire collected dataset and compare against analytical baselines.
-Predictions across unseen objects for each method are plotted in Fig.
-9.
-A traditional Hooke’s Law method without tactile sensing fails to distinguish between material types.
-Simple elastic and Hertzian analytical models exhibit stronger performance, but remain less than 30% accurate.
-Our hybrid learned model showcases dramatic improvements over all methods.
-In Table 1, performance is compared between different estimation methods and inputs.
-Performance metrics are computed for grasps of both seen and unseen objects.
-We evaluate models on a logarithmic scale given the range of Young’s moduli in our dataset.
-We compute error as the average difference between prediction and ground truth on a log10 scale.
-Predictions are considered accurate within the same order of magnitude of ground truth labels.
-This tolerance is deemed sufficient for downstream tasks of produce sorting based on the change in compliance of selected fruits and vegetables during ripening [54, 55, 56, 57].
-Figure 9: Young’s modulus predictions by material.
-Results for learned and analytical estimation methods are plotted, using predictions from 80 randomly sampled grasps per material category.
-Predictions in the gray region are considered sufficiently accurate.
-From the ablation study, we observe that our hybrid model drastically out-performs purely analytical and learning-based methods.
-RGB tactile images with markers are shown to be most favorable for this estimation task.
-There is not significant difference in results between models which consider force $F$ and width $w$ measurements.
-This may be due to the fact that all objects were grasped with identical peak force.
-Moreover, force is implicitly apparent through tactile difference images [12].
-Importantly, analytical estimates $\hat{E}_{\text{elastic}},\hat{E}_{\text{hertz}}$ are shown to improve prediction accuracy by nearly 10%.
-This indicates that a hybrid approach to compliance estimation is favorable over purely learning-based methods, which have been previously explored [14, 15, 1, 2, 3, 4].
-Prediction accuracy across shape and material categories for each method is investigated in Appendix E.
-We observe that our model performs independent of shape and can generalize well across materials.
+(b) Log10 error and accuracy are plotted across dataset categories of shape and material.
+Neural network results are from top-10 trained models evaluated on unseen objects.
 Inputs Seen Objects Unseen Objects
@@
 Depth $\times$ $\times$ 1.08 60.1% 1.04 59.4%
-Table 1: Estimation Architecture Ablation Study
+TABLE III: Estimation Architecture Ablation Study
+Furthermore, we utilize this reduced-domain training configuration to evaluate our architecture against previously-developed Shore hardness estimation models.
+We can only compare between predictions in this domain because previously-developed models have only been trained for use with rubber objects.
+Converting our model’s predictions from Young’s modulus to Shore 00 hardness [43], we find that our model reduces root mean square error (RMSE) compared to Yuan et al.
+[12] from 10.3 to 5.9 over simple-shaped objects and from 18.2 to 14.3 over arbitrarily-shaped objects, in units of Shore hardness.
+This observed improvement is likely due to our method’s integration of analytical estimates alongside tactile image-based CNN models, whereas previously methods were fully data-driven.
+Subsequently, we trained our neural network over the entire collected dataset and compare against analytical baselines.
+Predictions across unseen objects for each method are plotted in Fig.
+8a.
+A traditional Hooke’s Law method without tactile sensing fails to distinguish between material types.
+Simple elastic and Hertzian models exhibit stronger performance, but remain less than 30% accurate.
+Our hybrid learned model showcases dramatic improvements over all methods.
+Errors across all models may be partially attributed to potential inaccuracies in our dataset labelling process.
+In Table III, performance is compared between different ablations of our estimation methods and inputs.
+Performance metrics were computed for grasps of both seen and unseen objects.
+We evaluated models on a logarithmic scale given the range of Young’s moduli in our dataset.
+We computed error as the average difference between prediction and ground truth on a log10 scale.
+Accuracy was computed based on the percentage of predictions within the same order of magnitude of ground truth labels.
+From the ablation study, we observe that our hybrid model out-performs purely analytical and learning-based methods.
+RGB tactile images with markers are shown to be most favorable for this estimation task.
+There is only minor differences in results between models which consider normal force $F$ and width $w$.
+This may be due to the fact that all objects were grasped with identical peak force.
+Moreover, force is implicitly apparent through tactile difference images [8].
+Importantly, consideration of analytical estimates $\hat{E}_{\text{elastic}},\hat{E}_{\text{hertz}}$ is shown to improve prediction accuracy by nearly 10%.
+This indicates our hybrid approach to compliance estimation is favorable over purely learning-based methods, which have been previously explored [12, 14, 15, 29, 30, 31].
+As shown in Fig.
+8b, our model performs independent of shape and can generalize well across materials.
 Limitations: Our estimation system struggled to express precise differences in Young’s modulus between harder objects that stray further from the modulus of our sensor (i.e.
@@
 These results are expected given that harder objects are rigid relative to our soft sensor.
-The resolution of our tactile sensor is not high enough to detect minute differences in deformation between this range of compliances.
-Considering this fact, our estimation system exhibits only marginal effective improvement relative to previously developed Shore hardness estimation models [14].
-Further work must be done to create a truly universal compliance estimation system.
-7 Conclusion
+Moreover, the resolution of our tactile sensor is not high enough to detect minute differences in deformation between this range of compliances.
+Future experiments could be performed with harder sensor gels to verify if this corresponds in an increased ability to sense the compliance of harder objects.
+Additionally, alternative analytical models that preport to better model the non-linear, volume-conserving behavior of silicone could be explored.
+However, this limitation may represent an intrinsic bottleneck in the precision of depth-tracking in camera-based tactile sensors.
+VII Conclusion
 We have developed a novel system for estimating the Young’s modulus of unknown contacted objects using analytical and data-driven methods.
-This system can create estimates online from only a single parallel grasp using camera-based tactile sensors.
+This system can create online estimates from only a single parallel grasp using camera-based tactile sensors.
 It has been shown to perform independent of object shape and remains robust across material types.
-It can precisely distinguish between compliances of soft objects, but struggles to expressively differentiate between that of harder objects.
-The estimation architecture introduced in this paper could be applied for use with other camera-based tactile sensors capable of tracking depth.
-With more diverse tactile data, the accuracy of the system could be further improved.
-Given an increased ability to generalize, this compliance estimation system could be deployed alongside an intelligent high-level controller to inform decision-making in robotic manipulation scenarios by characterizing the compliance of unknown objects.
-This is particularly valuable when visually similar objects may have distinct material composition.
-As one example, it could be utilized to identify the ripeness of fruits or vegetables [1, 2, 4].
+It can precisely distinguish between compliances of soft objects and is more accurate over this material domain than the current state-of-the-art [12].
+Moreover, by parametrizing compliance via Young’s modulus, we are able to produce estimates on an absolute scale over many material categories.
+This generality is demonstrated through our diverse physical dataset.
+Given an ability to reliably produce estimations of compliance on a universal scale across different material regimes, this estimation system could be used alongside haptic displays to enhance teleoperation interfaces [3].
+With better tactile feedback to users, more performative policies for general robotic tasks and remote surgical operation could be developed.
+The estimation system has already been successfully deployed to measure compliance in real-time [44].
 ACKNOWLEDGMENTS
-The authors would like to thank Sandra Q. Liu, Jialiang (Alan) Zhao, Megha Tippur, Yuxiang Ma, and Laurence Willemet for their guidance and insight provided throughout the project’s duration.
-Appendix A Method of Dimensionality Reduction (MDR)
-As discussed in Section 3.2, we model contact between an unknown object and our flat, camera-based tactile sensor as a collision between an elastic sphere and an elastic half-space using Hertzian contact theory.
-Now, we will walk through the derivation for a constitutive relation using the method of dimensionality reduction (MDR) [58].
-A mechanical diagram of contact at time $t$ with labeled variables is provided in Fig.
-10.
-Figure 10: Method of dimensionality reduction (MDR) diagram.
-An elastic sphere is colliding with an elasic half-space at time $t$.
-Variables of contact are labeled.
-The two-dimensional radial stress field $p$ is transformed into a one-dimensional force field $q_{\text{1D}}$.
-This transformation can be observed through comparison of top and side views.
-First, we compute the maximum pressure of contact $p_{0}$ between both bodies in Eq.
-6.
-This is a simple expression of measured force $F$ and contact radius $a$.
-We can substitute $a$ to create an expression dependent on the aggregate elastic compliance of the bodies $E^{*}$ [47].
-$p_{0}(t)=\frac{3F(t)}{2\pi a^{2}(t)}=\frac{1}{\pi}{\left(\frac{F(t){E^{*}}^{2}}{R^{2}}\right)}^{\frac{1}{3}}$ (6)
-Using this maximum pressure $p_{0}$ at timestep $t$, we create the radial pressure field of contact $p(r,t)$. This expression comes directly from Hertzian contact theory [47].
-$p(r,t)=p_{0}(t)\sqrt{1-\frac{r^{2}}{a^{2}(t)}}$ (7)
-Now, we will apply the method of dimensionality reduction (MDR).
-This technique is used to solve the stress field for surface deformation based on the assumption that contact geometries are axisymmetric.
-By utilizing this method, we are able to reconstruct the deformation of the sensor surface.
-We begin by transforming coordinates of the pressure field into one-dimensional force density $q_{\text{1D}}$ through Eq.
-8.
-[…]
-From here, we apply a constitutive relationship on the sensor gel pad to get a one-dimensional displacement of the gel surface $u_{\text{1D}}$.
-This displacement is defined in the normal direction $\hat{z}$.
-By transforming to a single dimension, we restrict the displacement geometry to be axisymmetric about $\hat{z}$ from the center $x=0$.
-$u_{\text{1D}}(x,t)=\left(\frac{1-\nu_{\text{sensor}}^{2}}{E_{\text{sensor}}}\right)q_{\text{1D}}(x,t)$ (9)
-The maximum displacement will occur at $x=0$. We define this to be equivalent to our maximum measured depth $d$.
-$u_{\text{1D}}(0,t)=d(t)=\left(\frac{1-\nu_{\text{sensor}}^{2}}{E_{\text{sensor}}}\right)q_{\text{1D}}(0,t)$ (10)
-Then, by plugging in $x=0$, the expression for $q_{\text{1D}}$ simplifies considerably.
-[…]
-We solve this integral to get Eq. 13 which expresses maximum observed surface displacement $d$ in terms of normal force $F$.
-[…]
-$u_{\text{1D}}(0,t)=d(t)=\frac{1-\nu_{\text{sensor}}^{2}}{E_{\text{sensor}}}{\left(\frac{3{E^{*}}^{2}F(t)}{32R^{2}(t)}\right)}^{\frac{1}{3}}a(t)$ (13)
-We apply this relation for every timestep $t$ over the duration of a grasp.
-Using all measurements of $d$, $a$, and $F$, we can apply a linear least-squares algorithm to retrieve a best fit value $\hat{E}^{*}$.
-Then, this estimate is used with known mechanical properties of our sensor’s silicone gel pads to compute the Young’s modulus of the unknown grasped object $\hat{E}_{\text{hertz}}$, as shown in Eq.
-5.
-Without the MDR, we could not have solved directly for the deformation of the gel surface, meaning we could not create an expression relating $F$ directly to $d$ in terms of Young’s modulus.
-Instead, we could only derive a relation for the apparent deformation of both bodies with respect to each other.
-This apparent deformation is equal to the displacement between both bodies, or change in gripper width $\Delta w$ in our case.
-This is a classic tabulated solution for Hertzian contact between a sphere and a plane [47], written as Eq.
-14.
-$F(t)=2E^{*}a(t)\Delta w(t)$ (14)
-Notably, this model would not make use of tactile depth $d$ and relies solely on gripper width $w$ measurements.
-These gripper measurements exhibit a larger observed error than tactile depth measurements, which are higher-resolution and demonstrate sub-millimeter accuracy.
-For these reasons, we opt to use the MDR-derived solution.
-Appendix B Markers
-Figure 11: Markers. Tactile images from GelSight Wedge’s with and without markers. Tactile images are size 35mm by 25mm.
-Markers are small black dots painted in an array on the surface of the GelSight gel pad.
-As the surface of the gel deforms and stretches, these markers move.
-Thus, we can track their relative position to map the displacement of the gel pad surface in both shear and normal directions.
-Markers have been shown to be effective in estimating shear and normal forces from tactile images [12, 17].
-However, accurate force estimations require assumptions on contact geometry.
-In this work, we seek to estimate the compliance of unknown objects.
-Since we cannot make accurate assumptions on geometry, we measure forces through an external force gauge rather than through our GelSight sensors.
-In any case, we can use markers to increase the amount of dynamic information encoded in a tactile image.
-Without them, we could not perceive shear displacement.
-In our experiment, we have created one GelSight Wedge [17] with markers painted on the surface.
-This was done to investigate how beneficial markers may be for estimating a dynamic property of contact like Young’s modulus.
-The difference between tactile images with or without markers for two objects is displayed in Fig.
-11.
-These markers help track surface displacement, but they may occlude fine features of texture, like screw threads.
-Appendix C Processing Grasp Data
-We collect data over a set number of automated parallel grasps for every object in our dataset.
-Normal force $F$, width $w$, and maximum depth $d$ are plotted for an example grasp of a rigid sphere in Fig.
-12a.
-Grasps are executed at nearly constant velocity to a peak force of approximately 30 N.
-These measurements are clipped to the loading sequence based on force thresholding.
-Once force passes an empirically-defined threshold of 0.75 N, we consider the object to be in-contact with the GelSight sensor and define this as timestep $t=0$.
-We clip data from this point until the peak recorded force is reached.
-As expected in an elastic regime, we observe that force and depth measurements are high correlated from both GelSight sensors.
-Raw collected data normalized per grasp is plotted in Fig.
-12b.
-Measurements are normalized according to their peak value over each respective grasp.
-These measurements are randomly sampled from grasps across the entire dataset.
-All measurements outside of 25% correlation bounds are discarded.
-Over 90% of measured points lie within these bounds for both sensors.
-We observe that there is no distinguishable difference between measurements collected from one sensor to the other.
-Figure 12: Grasping data.
-(a) Grasp data is plotted for an example grasp of a rigid sphere.
-The chosen loading sequence from initial contact to peak force is highlighted in gray.
-(b) Normalized force and depth measurements from all collected grasps are plotted.
-The correlation coefficient between these measurements is computed as 0.84 and bounds are placed at 25% error mark.
-Appendix D Comparison Against Shore 00 Hardness Baseline
-Previous works have developed learning-based methods to estimate Shore hardness using camera-based tactile sensors without the incorporation of analytical modelling [40, 14, 5, 15, 1].
-Notably, Yuan et al.
-introduced an architecture to estimate Shore 00 hardness of various sample geometries using tactile images captured from a GelSight sensor [14].
-To demonstrate the performance of our novel hybrid estimation system, we directly compare it against results from Yuan et al.
-As discussed in Section 6, we can constrain our dataset to predict compliance for only rubber objects.
-There are 82 rubber objects in our dataset of 285 total objects.
-The compliance of these rubber objects aligns with the Shore 00 hardness scale, which spans approximately from a Young’s modulus of 6 kPa to 2 MPa.
-We can directly evaluate and compare predictions of Young’s modulus with Shore 00 hardness by utilizing established models that convert between the two measures [53].
-Using the rubber subset of our dataset, we replicate experiments conducted by Yuan et al.
-and compare their quoted results against our own [14].
-We use RGB tactile images with markers and analytical estimates as the input to our model.
-First, we train an instance of our hybrid estimation system to predict compliance using only simple shapes such as spheres, cylinders, and rectangular objects.
-This model is validated across seen shapes with unseen compliances.
-Predictions from our model are subsequently converted to Shore 00 hardness to calculate error over the validation set.
-Second, we train another instance of our hybrid model and validate it across unseen objects with arbitrary or irregular shapes.
-It is important to note that Yuan et al.
-evaluate over hand-collected data for the latter experiment, whereas all of our data collection is performed on robot hardware.
-Results from our experiments are presented in Table 2, where root mean-squared error (RMSE) is reported in units of Shore 00 hardness.
-Method Dataset $R^{2}$ RMSE
-Shore 00 Hardness Simple Shapes 0.85 10.3
-Model [14] Arbitrary Shapes 0.39 18.2
-Top-1 Hybrid NN Simple Shapes 0.69 5.9
-(Our Method) Arbitrary Shapes 0.37 14.3
-Table 2: Shore 00 Hardness Prediction Error Comparison
-Our estimation system exhibits decreased error in both experiments across both simple and complex shapes, though the margin of improvement is modest.
-These results suggest that our novel hybrid estimation architecture offers some performance enhancement over previously developed learning-based approaches for predicting the compliance of soft rubbers.
-We can likely attribute this decrease in error to the incorporation of analytical modeling.
-Appendix E Dataset Diversity & Performance
-The diversity of the collected dataset is characterized through categories of material and shape in Tables 4 and 4.
-As intended, the most common material of objects in the dataset is rubber, which includes 28.6% of the objects in the dataset.
-This is done to bias the dataset toward the Young’s modulus of our sensor’s gel pads.
-There are significant percentages of metal and plastic objects, making up 25.1% and 24.0% of the dataset respectively.
-Over 50% of objects in the dataset have irregular shapes.
-This is done to diversify training data in an attempt and learn Young’s modulus independent of object shape or contact profile.
-Still, there are substantial percentages of simple shapes for performance evaluation.
-Table 3: Dataset Material Distributions
-Material Percent (%)
-Rubber 28.6%
-Metal 25.1%
-Plastic 24.0%
-Wood 8.4%
-Foam 6.3%
-Other 7.6%
-Table 4: Dataset Shape Distributions
-Shape Percent (%)
-Cylinder 20.2%
-Sphere 11.5%
-Rectangular 15.3%
-Hexagonal 2.1%
-Irregular 50.9%
-Results of investigated methods on unseen objects across shape and material categories are compared in Fig.
-13.
-We observe that neural network methods give the lowest log10 error and accuracy in nearly every category.
-Moreover, the accuracy increases and error decreases with the complexity of analytical models.
-Our Hertzian method performs best over spheres, which is expected given that this model assumes spherical contact geometry.
-Most critically, it is shown that our trained model can learn Young’s modulus independent of shape, given that prediction accuracy and error remain consistent across different object shapes.
-The model performs robustly across different categories of materials as well.
-Figure 13: Estimation methods performance.
-Log10 error and accuracy are plotted for each estimation method across dataset categories of shape and material.
-Neural network results are from top-10 trained models evaluated on unseen objects.
+The authors would like to thank Megha Tippur, Sandra Q. Liu, and Yuxiang Ma for their valuable insights and guidance.

### v4

@@
 Importantly, this estimation system performs irrespective of object geometry and demonstrates increased robustness across material types.
+Code is available on GitHub and collected data is available on HuggingFace.
 I Introduction
@@
 By incorporating analytical models, we improve generalization across materials, while using a learned CNN to account for geometric complexities.
-We choose to label compliance as Young’s modulus to align with analytical models and so that our system produces estimations on a universal scale.
+We label compliance as Young’s modulus to align with analytical models and produce estimations on a universal scale.
 In this work, we present the following contributions:
@@
 This model assumes symmetry between parallel fingers.
-We define $t=0$ to be the moment of first contact, where $F(t=0)$ must be greater than 0.75 N.
+We define $t=0$ to be the moment of first contact, where $F(t=0)$ must be greater than an impirically chosen threshold of 0.75 N.
 $\sigma(t)=\frac{F(t)}{A(t)}$ (1)
@@
 Unlike previous methods, Hertzian mechanics closely analyzes geometry through contact.
-It has poreviously been applied to model soft contact in robotics with camera-based tactile sensors [23, 25, 22, 13].
+It has been applied to model soft contact in robotics with camera-based tactile sensors [23, 25, 22, 13].
 GelSight sensors are made with elastic silicone rubber [16].
@@
 All displayed tactile images are size 35mm by 25mm.
-V Data Collection & Training
+V Experiments
 V-A Hardware Setup
@@
 [12] from 10.3 to 5.9 over simple-shaped objects and from 18.2 to 14.3 over arbitrarily-shaped objects, in units of Shore hardness.
-This observed improvement is likely due to our method’s integration of analytical estimates alongside tactile image-based CNN models, whereas previously methods were fully data-driven.
+This observed improvement is likely due to our method’s integration of analytical estimates alongside tactile image-based CNN models, whereas previous methods were fully data-driven.
 Subsequently, we trained our neural network over the entire collected dataset and compare against analytical baselines.